### v2

# Probabilistic Graph Cuts

Abstract
Probabilistic relaxations of graph cuts offer a differentiable alternative to spectral clustering, enabling end-to-end and online learning without eigendecompositions, yet prior work centered on RatioCut and lacked general guarantees and principled gradients. We present a unified probabilistic framework that covers a wide class of cuts, including Normalized Cut. Our framework provides tight analytic upper bounds on expected discrete cuts via integral representations and Gauss hypergeometric functions with closed-form forward and backward. Together, these results deliver a rigorous, numerically stable foundation for scalable, differentiable graph partitioning covering a wide range of clustering and contrastive learning objectives.
Probabilistic Graph Cuts
Ayoub Ghriss
Department of Computer Science
University of Colorado, Boulder
ayoub.ghriss@colorado.edu
1 Introduction
Self-supervised learning (SSL) is the backbone of modern representation learning at scale across vision, speech, and language. Large-scale pretraining coupled with lightweight fine-tuning or even zero-shot transfer closes the gap to fully supervised baselines. For example, in vision-language, CLIP-style contrastive pretraining delivers robust zero-shot transfer across several benchmarks (Radford et al., 2021). In speech, wav2vec 2.0 enables competitive recognition with orders of magnitude fewer labeled hours than conventional pipelines (Baevski et al., 2020).
The majority of successful SSL objectives fall into two families. Contrastive methods maximize agreement between augmented positives while separating negatives, typically via InfoNCE (van den Oord et al., 2018). Canonical instances include SimCLR (Chen et al., 2020), MoCo (He et al., 2020), SwAV (Caron et al., 2020), and theory framing performance in terms of alignment and uniformity on the sphere (Wang & Isola, 2020). A second line removes explicit negatives with non-contrastive joint-embedding objectives such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), Barlow Twins (Zbontar et al., 2021), VICReg (Bardes et al., 2022), and self-distillation (DINO) (Siméoni et al., 2025). Orthogonal to these, masked prediction pretrains encoders by reconstructing masked content: BERT for language (Devlin et al., 2019), BERT for Images (BEiT), Masked Autoencoders (MAE) for images (Bao et al., 2022; He et al., 2021), and HuBERT/wav2vec 2.0 for speech (Hsu et al., 2021). Finally, multimodal contrastive pretraining (ALIGN) scales this recipe to billions of image–text pairs for broad zero-shot capabilities (Jia et al., 2021).
A parallel stream connects SSL with clustering and partitioning: DeepCluster iterates k-means assignments and network updates (Caron et al., 2018); Invariant Information Clustering (IIC) maximizes mutual information between paired views’ discrete assignments (Ji et al., 2019); Self-Labeling (SeLa) and SwAV exploit online clustering and optimal transport (YM. et al., 2020). Yet, despite this progress, widely used partitioning losses rooted in classical graph cuts; RatioCut and Normalized Cut (Ncut) remain challenging to optimize end-to-end. Their standard spectral relaxations are based on eigendecompositions (Hagen & Kahng, 1992; Shi & Malik, 2000), which complicate large-scale differentiation and online training.
Contrastive and masked objectives are typically pairwise or reconstruction surrogates; in contrast, cut-based criteria operate at the partition level, directly penalizing cross-cluster affinity relative to within-cluster volume. Our results supply a tractable bridge: they turn the discrete objectives into smooth surrogates with guarantees, compatible with modern encoders and large-batch training. we formalize conditions under which common SSL pipelines (e.g., SimCLR/CLIP-style graphs) emerge as special cases of our envelope when the similarity graph is built from the current embeddings, linking alignment–uniformity improvements to monotone decreases of our upper bound.
In this paper, we provide a unified and differentiable probabilistic relaxation of graph cuts (including Ncut) with closed-form forward/backward via hypergeometric polynomials, yielding a stable, scalable surrogate that avoids eigendecompositions. We prove monotonicity, separate convexity, and data-dependent Lipschitz/curvature bounds for principled optimization; give two-sided, zero-aware AM–GM gap control with a simple penalty; establish minibatch concentration; and handle heterogeneous degrees through a Hölder-product binning scheme with a colinearity-guided selection rule. The resulting objective upper-bounds the expected cut and tightens during training, and common SSL setups (SimCLR/CLIP) emerge as special cases by building the graph from current embeddings. Taken together, the theory, algorithms, and plug-and-play implementation form a coherent, rigorously justified approach to scalable, end-to-end graph partitioning.
2 Preliminaries
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\bm{W}})$ be an undirected weighted graph on $n=|{\mathcal{V}}|$ vertices with a symmetric, elementwise nonnegative adjacency matrix ${\bm{W}}\in\mathbb{R}^{n\times n}$; assume ${\bm{W}}_{ii}=0$. Define the degree of $i\in{\mathcal{V}}$ by $d_{i}\vcentcolon=\sum_{j\in{\mathcal{V}}}{\bm{W}}_{ij}$ and the degree matrix ${\bm{D}}\vcentcolon=\mathop{\mathrm{diag}}(d_{1},\ldots,d_{n})$.
For ${\mathbb{A}}\subseteq{\mathcal{V}}$, let $\overline{{\mathbb{A}}}\vcentcolon={\mathcal{V}}\setminus{\mathbb{A}}$ and identify $\mathbb{A}$} by its indicator vector ${\bm{1}}_{{\mathbb{A}}}\in\{0,1\}^{n}$. The cut associated with $\mathbb{A}$} is:
$\operatorname*{cut}({\mathbb{A}},\overline{{\mathbb{A}}})\;\vcentcolon=\sum_{(v_{i},v_{j})\in{\mathbb{A}}\times\overline{{\mathbb{A}}}}{\bm{W}}_{ij}\;=\;{\bm{1}}_{{\mathbb{A}}}^{\top}{\bm{W}}\,{\bm{1}}_{\overline{{\mathbb{A}}}},$
and the associated volume-normalized cut is
$\operatorname*{VolCut}({\mathbb{A}})\;\vcentcolon=\;\frac{\operatorname*{cut}({\mathbb{A}},\overline{{\mathbb{A}}})}{\operatorname*{vol}({\mathbb{A}})},$ (1)
where the volume is $\operatorname*{vol}({\mathbb{A}})\vcentcolon=\sum_{v_{i}\in{\mathbb{A}}}s(v_{i})$ for a given vertex weight function $s:{\mathcal{V}}\to\mathbb{R}_{>0}$. For example, the ratio cut uses $s(v_{i})\equiv 1$ so $\operatorname*{vol}\left({\mathbb{A}}\right)\equiv|{\mathbb{A}}|$, whereas the normalized cut uses $s(v_{i})=d_{i}$.
Fix $s$ and write $s_{i}\vcentcolon=s(v_{i})$. The goal is to find a $k$-way partition ${\mathcal{C}}_{k}=\{{\mathbb{C}}_{\ell}\}_{\ell=1}^{k}$ of $\mathcal{V}$} that minimizes the volume-normalized graph cut:
$\text{GraphCut}({\mathcal{C}}_{k})\;\vcentcolon=\;\frac{1}{2}\sum_{\ell=1}^{k}\operatorname*{VolCut}({\mathbb{C}}_{\ell}).$ (2)
2.1 Probabilistic Relaxation
The Probabilistic Ratio-Cut (PRCut) (Ghriss & Monteleoni, 2025) adopts a probabilistic relaxation of $k$-way clustering. Let ${\mathbf{a}}_{\ell}\in\{0,1\}^{n}$ be the random indicator of the cluster $\mathbb{C}$$_{\ell}$. The clustering ${\mathcal{C}}_{k}$ is parameterized by a row-stochastic matrix ${\bm{P}}\in[0,1]^{n\times k}$ with $\sum_{\ell=1}^{k}{\bm{P}}_{i\ell}=1$ and ${\bm{P}}_{i\ell}=\operatorname{Pr}\left({\textnormal{a}}_{\ell,i}=1\right)=\operatorname{Pr}\left(v_{i}\in{\mathbb{C}}_{\ell}\right)$.
The expected graph cut is defined as:
$\text{GraphCut}({\bm{P}})\;\vcentcolon=\;\frac{1}{2}\sum_{\ell=1}^{k}\mathbb{E}\left[\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\right],$ (3)
where
$\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\;\vcentcolon=\;\frac{{\mathbf{a}}_{\ell}^{\top}{\bm{W}}\,({\bm{1}}-{\mathbf{a}}_{\ell})}{\sum_{i=1}^{n}s_{i}\,{\textnormal{a}}_{\ell,i}}.$ (4)
The following bound underpins the PRCut framework:
Proposition 2.1 (PRCut bound (Ghriss & Monteleoni, 2025)).
For the ratio cut ($s_{i}\equiv 1$ for all $i\in\{1,\ldots,n\}$):
$\mathbb{E}\left[\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\right]\leq\frac{1}{n\,\overline{{\bm{P}}}_{:,\ell}}\sum_{i,j=1}^{n}{\bm{W}}_{ij}\bigl({\bm{P}}_{i\ell}+{\bm{P}}_{j\ell}-2\,{\bm{P}}_{i\ell}{\bm{P}}_{j\ell}\bigr),$
where $\overline{{\bm{P}}}_{:,\ell}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}{\bm{P}}_{i\ell}$ denotes the expected fraction of vertices assigned to $\mathbb{C}$$_{\ell}$.
In this paper, we derive tighter, more general bounds for an arbitrary vertex-weight function $s$, with concentration guarantees and gradients that are compatible with first-order optimization.
3 Proposed methods
By symmetry in Equation 3, it suffices to bound the expected $\operatorname*{VolCut}$ for a single cluster. Fix a cluster $\mathbb{C}$} and drop the index $\ell$. Let ${\mathbf{a}}\in\{0,1\}^{n}$ be its random indicator with independent coordinates ${\mathbf{a}}_{i}\sim\mathrm{Bernoulli}(p_{i})$, and let ${\bm{p}}=(p_{1},\ldots,p_{n})^{\top}\in[0,1]^{n}$.
$\widehat{\operatorname*{VolCut}}({\mathbf{a}})\;=\;\sum_{i,j=1}^{n}{\bm{W}}_{ij}\,\frac{{\mathbf{a}}_{i}(1-{\mathbf{a}}_{j})}{\sum_{l=1}^{n}s_{l}\,{\mathbf{a}}_{l}}.$ (5)
Consider the $(i,j)=(1,2)$ term (note that ${\bm{W}}_{11}={\bm{W}}_{22}=0$):
$\mathbb{E}\left[\frac{{\mathbf{a}}_{1}(1-{\mathbf{a}}_{2})}{\sum_{i=1}^{n}s_{i}\,{\mathbf{a}}_{i}}\right]\;=\;p_{1}(1-p_{2})\;\mathbb{E}\left[\frac{1}{\,s_{1}+\sum_{i=3}^{n}s_{i}\,{\mathbf{a}}_{i}\,}\right].$
Thus, we must evaluate expectations of the form $\mathbb{E}\left[1/(q+{\textnormal{x}})\right]$ with $q>0$ and ${\textnormal{x}}=\sum_{l=3}^{n}s_{l}\,{\mathbf{a}}_{l}$. It turns out that x follows a generalized Poisson–Binomial distribution. The Poisson–Binomial distribution is well studied and has applications across seemingly unrelated areas (Chen & Liu, 1997; Cam, 1960). We use its generalized form:
Definition 3.1 (Generalized Poisson–Binomial (GPB)).
Let ${\bm{\alpha}}\in[0,1]^{m}$ and $\theta_{i}<\beta_{i}$ be real constants, and let ${\textnormal{r}}_{i}\sim\mathrm{Bernoulli}(\alpha_{i})$ independently. The random variable ${\textnormal{x}}\;=\;\sum_{i=1}^{m}\big(\theta_{i}(1-{\textnormal{r}}_{i})+\beta_{i}{\textnormal{r}}_{i}\big)$ follows a generalized Poisson–Binomial distribution (Zhang et al., 2017).
In our setting, $m\vcentcolon=n-2$, ${\bm{\alpha}}=(p_{3},\ldots,p_{n})$, and the weights are $\theta_{i}=0$ and $\beta_{i}=s_{i}$, so ${\textnormal{x}}=\sum_{i=3}^{n}s_{i}\,{\textnormal{r}}_{i}$. We denote this special case by $\operatorname{GPB}({\bm{\alpha}},{\bm{\beta}})$, and compute its probability generating function (PGF) $G_{\textnormal{x}}$:
$G_{\textnormal{x}}(t)\vcentcolon=\mathbb{E}\left[t^{{\textnormal{x}}}\right]=\prod_{i=1}^{m}\big(1-\alpha_{i}+\alpha_{i}\,t^{\,\beta_{i}}\big),\;t\in[0,1].$ (6)
The target expectation can now be computed via the identity $x^{-1}=\int_{0}^{1}t^{x-1}\,dt$ for $x>0$ (see Section A.2):
Lemma 3.2 (Integral representation).
For any $q>0$, we have:
$\mathbb{E}\left[\frac{1}{q+{\textnormal{x}}}\right]=\;\vcentcolon=\;\int_{0}^{1}t^{\,q-1}\,\prod_{i=1}^{m}\big(1-\alpha_{i}+\alpha_{i}\,t^{\,\beta_{i}}\big)\,dt.$ (7)
We denote this integral by ${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})$.
For the ratio cut, $q=1$ and $\beta_{i}\equiv 1$, and PRCut uses the bound $\mathbb{E}\left[\tfrac{1}{1+{\textnormal{x}}}\right]\leq(\sum_{i}\alpha_{i})^{-1}$. In this work, $s$ need not be constant, so different tools are required.
We first consider the case $\beta_{i}\equiv\beta$ and recall Gauss’s hypergeometric function $_{2}F_{1}$ (Chambers, 1992), defined for $|z|<1$ by the absolutely convergent power series:
$_{2}F_{1}(a,b;c;z)\;=\;\sum_{k=0}^{\infty}\frac{(a)_{k}\,(b)_{k}}{(c)_{k}}\,\frac{z^{k}}{k!},$ (8)
where $(x)_{k}\vcentcolon=x(x+1)\cdots(x+k-1)$ and $(x)_{0}\vcentcolon=1$ is the rising factorial.
Lemma 3.3 (Euler’s identity).
If $c>b>0$ and $z\in[0,1]$, then $_{2}F_{1}(a,b;c;z)$ is equal to:
$\frac{\Gamma(c)}{\Gamma(b)\,\Gamma(c-b)}\int_{0}^{1}t^{\,b-1}(1-t)^{\,c-b-1}\,(1-zt)^{-a}\,dt,$ (9)
where $\Gamma$ denotes the gamma function.
A useful identity is the derivative formula:
$\frac{d}{dz}\,{}_{2}F_{1}(a,b;c;z)\;=\;\frac{ab}{c}\,{}_{2}F_{1}(a+1,b+1;c+1;z),$ (10)
which, in particular, implies the following:
Lemma 3.4 (Properties of $_{2}F_{1}$).
Let $m\in\mathbb{N}$, $b>0$, and $c>b$. On $[0,1]$, the function $f(z)\vcentcolon={}_{2}F_{1}(-m,b;c;z)$ is a degree-$m$ polynomial that is decreasing, convex, and $L$-Lipschitz with $L=\tfrac{mb}{c}$.
The integral in Lemma 3.2 admits a computable and differentiable upper bound (proof in Section A.3).
Theorem 3.5 (Hypergeometric bound).
Assume $\beta_{i}\equiv\beta>0$. For any $q>0$,
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\;\leq\;\frac{1}{q}\;{}_{2}F_{1}\!\left(-m,\,1;\,\frac{q}{\beta}+1;\,\bar{\alpha}\right),$ (11)
where $\bar{\alpha}\vcentcolon={\frac{1}{m}}\sum_{i=1}^{m}\alpha_{i}$. We denote the right-hand side by ${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)$.
3.1 The AM-GM gap
To quantify the tightness of the bound from Theorem 3.5, we follow a similar structure for the pointwise Arithmetic Mean-Geometric Mean (AM-GM) gap.
Proposition 3.6 (Integrated AM–GM gap).
Let $\beta_{i}\equiv\beta>0$. Define $h(t)=t^{q-1}\left(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\right)^{m}$ and:
$\underline{\Delta}(q,{\bm{\alpha}})$ $\vcentcolon=\int_{0}^{1}h(t)\Bigl(1-e^{-\gamma(t)\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
$\overline{\Delta}(q,{\bm{\alpha}})$ $\vcentcolon=\int_{0}^{1}h(t)\Bigl(1-e^{-\theta(t)\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
with $\gamma(t)\vcentcolon=\tfrac{m}{2}(1-t^{\beta})^{2}$ and $\theta(t)\vcentcolon=\gamma(t)/t^{2\beta}$. Then:
$\underline{\Delta}(q,{\bm{\alpha}})\leq{\mathcal{H}}_{\beta}(q;\bar{\alpha},m)\;-\;{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\leq\overline{\Delta}(q,{\bm{\alpha}}),$ (12)
with equality throughout iff $\mathrm{Var}({\bm{\alpha}})=0$.
A convenient corollary gives an explicit upper bound.
Corollary 3.7 (Simple upper bound).
Under the conditions of Proposition 3.6, for any $q>0$,
${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)-{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\leq\frac{m}{2}\mathrm{Var}({\bm{\alpha}})\int_{0}^{1}h(t)\theta(t)\,dt.$
See Section A.4 for the proofs of Proposition 3.6 and its corollary.
Zero-aware gap control.
By Equation 7, coordinates with $\alpha_{i}=0$ contribute the factor $(1-\alpha_{i}+\alpha_{i}t^{\beta})\equiv 1$ and thus make no contribution to ${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})$. The AM–GM gap in Proposition 3.6 over-penalizes configurations with many inactive entries: zeros still inflate $\mathrm{Var}({\bm{\alpha}})$ even though they do not affect the product inside the integral. We therefore replace the plain variance by a zero-aware weighted dispersion that vanishes at $\alpha_{i}=0$. Let $\omega_{0}(x)\vcentcolon=x$ (more generally, $\omega_{0}(x)=x^{a},a\in[1,2]$), and define:
$\Omega\;\vcentcolon=\;\sum_{i=1}^{m}\omega_{0}(\alpha_{i}),\qquad\bar{\alpha}^{\omega_{0}}\;\vcentcolon=\;\frac{1}{\Omega}\sum_{i=1}^{m}\omega_{0}(\alpha_{i})\,\alpha_{i},$ (13)
$\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\vcentcolon=\;\frac{1}{\Omega}\sum_{i=1}^{m}\omega_{0}(\alpha_{i})\bigl(\alpha_{i}-\bar{\alpha}^{\omega_{0}}\bigr)^{2},$ (14)
with $\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})=0$ if $\Omega=0$.
Let ${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)$ be the envelope from Theorem 3.5. Define the second forward $\beta$-difference:
${\mathcal{A}}(q;\bar{\alpha},m)\;\vcentcolon=\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\;{\mathcal{H}}_{\beta}\!\bigl(q+r\beta;\bar{\alpha},m\bigr).$ (15)
and $\widetilde{{\mathcal{A}}}(q;\bar{\alpha},m)=\frac{\partial}{\partial\bar{\alpha}}{\mathcal{A}}(q;\bar{\alpha},m)$.
Proposition 3.8 (Zero-aware AM-GM gap).
Under the assumptions of Proposition 3.6 with common $\beta>0$, a zero-aware replacement for the simple upper bound of Corollary 3.7 is:
${\mathbb{A}}(q,{\bm{\alpha}},m)\vcentcolon=\frac{m}{2}\;\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr).$ (16)
In particular, coordinates with $\alpha_{i}\equiv 0$ incur zero penalty, while $\alpha_{i}=1$ retains full influence through $\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})$; when $\omega_{0}\!\equiv\!1$ we recover Corollary 3.7.
The main takeaway from Equation 16 is that the AM-GM gap can be reduced by pushing the active entries towards a common value. That is particularly true for the case where $\alpha_{i}\in\left\{0,1\right\}$ and validates the claim made by Ghriss & Monteleoni (2025) about the tightness of their PRCut upper-bound in the deterministic setting.
More details are provided in Section A.5 about the derivation of various quantities.
3.2 Concentration for minibatch estimators
Fix $m\in\mathbb{N}$, $q>0$, $\beta>0$. Let $c=\tfrac{q}{\beta}$ and $\{\alpha_{i}\}_{i=1}^{m}\subset[0,1]$ be the (fixed) population, with mean $\bar{\alpha}\!=\!m^{-1}\sum_{i}\alpha_{i}$ and population variance $\sigma^{2}\!\vcentcolon=\!m^{-1}\sum_{i}(\alpha_{i}-\bar{\alpha})^{2}$. Form a minibatch $S=(I_{1},\dots,I_{n})$ of size $n$ by sampling indices with replacement, and define the plug-in estimator:
$\hat{\alpha}_{S}\;\vcentcolon=\;\tfrac{1}{n}\sum_{r=1}^{n}\alpha_{I_{r}},\qquad\tilde{H}(S)\;\vcentcolon=\;{\mathcal{H}}_{\beta}\!\bigl(q;\hat{\alpha}_{S},m\bigr).$ (17)
By Lemma 3.4, ${\mathcal{H}}_{\beta}(q;\cdot,m)$ is decreasing, convex and $L$-Lipschitz with:
$L\;=\;\frac{m}{q\left(c+1\right)}.$ (18)
Differentiating equation 10 again gives, for all $z\in[0,1]$,
$\frac{d^{2}}{dz^{2}}\,{\mathcal{H}}_{\beta}(q;z,m)\;\leq\;\underbrace{\frac{1}{q}\cdot\frac{2m(m-1)}{c(c+1)}}_{\!\!\!\vcentcolon=K},$ (19)
Since ${\mathcal{H}}_{\beta}(q;\cdot,m)$ is convex, the second-order Taylor bound from Equation 19 yields:
$0\;\leq\;\mathbb{E}\,\tilde{H}(S)\;-\;{\mathcal{H}}_{\beta}\!\bigl(q;\bar{\alpha},m\bigr)\;\leq\;\frac{K}{2}\,\frac{\sigma^{2}}{n},$ (20)
where the variance identity holds for sampling with replacement.
Changing one element of $S$ changes $\hat{\bar{\alpha}}_{S}$ by at most $1/n$, hence by equation 18 the function $S\mapsto\tilde{H}(S)$ changes by at most $L/n$. McDiarmid’s inequality gives, for all $\varepsilon>0$,
$\Pr\!\left(\bigl|\tilde{H}(S)-\mathbb{E}\,\tilde{H}(S)\bigr|\geq\varepsilon\right)\;\leq\;2\exp\!\left(-\frac{2n\,\varepsilon^{2}}{L^{2}}\right).$ (21)
Combining equation 20 and equation 21 with a triangle inequality yields the following finite-sample guarantee.
Proposition 3.9 (Concentration of the minibatch envelope).
With probability at least $1-\delta$,
$\bigl|\tilde{H}(S)-{\mathcal{H}}_{\beta}(q;\bar{\alpha},m)\bigr|\leq L\,\sqrt{\frac{1}{2n}\log\!\frac{2}{\delta}}+\frac{K}{2}\,\frac{\sigma^{2}}{n},$ (22)
where $L$ and $K$ from Equations 18 and 19.
Proof.
By McDiarmid, with probability $\geq 1-\delta$, $|\tilde{H}(S)-\mathbb{E}\,\tilde{H}(S)|\leq L\sqrt{\tfrac{1}{2n}\log(2/\delta)}$. Add and subtract ${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)$ and use $0\leq\mathbb{E}\,\tilde{H}(S)-{\mathcal{H}}_{\beta}(q;\bar{\alpha},m)\leq\tfrac{K}{2}\mathrm{Var}(\hat{\bar{\alpha}}_{S})$ from equation 20, with $\mathrm{Var}(\hat{\bar{\alpha}}_{S})=\sigma^{2}/n$. ∎
3.3 Heterogeneous degrees
When $(\beta_{i})_{i}$ vary, directly using a single $\beta$ loses heterogeneity. We partition indices into $d$ disjoint bins $S_{1},\dots,S_{d}$ based on their $\beta_{i}$ values. Let $m_{j}\vcentcolon=|S_{j}|$, $\bar{\alpha}_{j}\vcentcolon=m_{j}^{-1}\sum_{i\in S_{j}}\alpha_{i}$, and define the bin interval $B_{j}=[b_{j-1},b_{j}]$ with representatives $\beta_{j}^{\star}\in B_{j}$ specified below.
For $t\in[0,1]$ and $\alpha\in[0,1]$, the map $\beta\mapsto(1-\alpha+\alpha\,t^{\beta})$ is nonincreasing. Hence, for any fixed bin $S_{j}$ and any choice $\beta_{j}^{\star}\leq\beta_{i}$ for all $i\in S_{j}$, we have:
$\prod_{i\in S_{j}}\Bigl(1-\alpha_{i}+\alpha_{i}\,t^{\beta_{i}}\Bigr)\;\leq\;\prod_{i\in S_{j}}\Bigl(1-\alpha_{i}+\alpha_{i}\,t^{\beta_{j}^{\star}}\Bigr).$ (23)
Applying Jensen in $\alpha$ to the RHS (log is concave in $\alpha$ for fixed $t^{\beta}$) gives, for each $j$,
$\prod_{i\in S_{j}}\Bigl(1-\alpha_{i}+\alpha_{i}\,t^{\beta_{j}^{\star}}\Bigr)\;\leq\;\Bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}\,t^{\beta_{j}^{\star}}\Bigr)^{m_{j}}.$ (24)
Let $m=\sum_{j=1}^{d}m_{j}$ and recall the envelope ${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)$ from Theorem 3.5. We now control the heterogeneous case.
Theorem 3.10 (Binned Hölder bound).
Let $q>0$ and partition $\{1,\ldots,m\}$ into bins $S_{1},\dots,S_{d}$. Choose representatives $\beta_{j}^{\star}\in B_{j}$ satisfying $\beta_{j}^{\star}\leq\beta_{i}$ for every $i\in S_{j}$ (e.g., left endpoints). Then
${\mathcal{I}}\bigl(q;{\bm{\alpha}},{\bm{\beta}}\bigr)\;\;\leq\;\;\prod_{j=1}^{d}\Bigl[\;{\mathcal{H}}_{\beta_{j}^{\star}}\!\bigl(q;\bar{\alpha}_{j},m\bigr)\;\Bigr]^{\frac{m_{j}}{m}}.$ (25)
Hölder inequality is tight iff the functions $\{f_{j}\}_{j=1}^{d}$ are pairwise proportional (colinear) in $L^{p_{j}}$: there exist constants $\kappa_{j}>0$ and a common shape $\phi$ such that $f_{j}(t)=\kappa_{j}\,\phi(t)$ for almost every $t\in[0,1]$. In our construction,
$f_{j}(t)\propto t^{\frac{q-1}{p_{j}}}\Bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}\,t^{\beta_{j}^{\star}}\Bigr)^{m}.$
Hence near-tightness is promoted when, across bins, the curves $t\mapsto(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}\,t^{\beta_{j}^{\star}})$ have similar shapes, and, within bins, replacing $\beta_{i}$ by $\beta_{j}^{\star}$ induces minimal distortion.
3.4 Optimization objective
We now put everything together to define the optimization objective of our probabilistic graph cut framework. For cluster $\ell$, the expected contribution of edge $(i,j)$ is:
$\frac{1}{s_{i}}{\bm{W}}_{ij}\,{\bm{P}}_{i\ell}(1-{\bm{P}}_{j\ell}){\mathcal{I}}\!\bigl(s_{i};{\bm{P}}_{-\{i\ell\}},{\bm{s}}_{-\{il\}}\bigr)$
Fix $\ell\in\{1,\dots,k\}$ and partition indices into $d$ bins $S_{\ell 1},\ldots,S_{\ell d}$ by their exponents $\beta_{u}\equiv s_{u}$ (e.g., degree-based); let $m_{\ell j}\vcentcolon=|S_{\ell j}|$, $m_{\ell}\vcentcolon=\sum_{j}m_{\ell j}$, and
$\bar{p}_{\ell j}\;\vcentcolon=\;\frac{1}{m_{\ell j}}\sum_{u\in S_{\ell j}}{\bm{P}}_{u\ell},\qquad w_{\ell j}\;\vcentcolon=\;\frac{m_{\ell j}}{m_{\ell}}.$
Choose representatives $\beta_{\ell j}^{\star}\leq s_{u}$ for all $u\in S_{\ell j}$ (e.g., the bin’s left endpoint or in–bin minimum) so that the bound direction is preserved (Section 3.3).
For a fixed $q>0$, Theorem 3.10 yields the per–cluster integrand bound. Plugging $q=s_{i}$ for each source vertex $i$ gives the per-vertex envelope:
$\Phi_{\ell}(q)\;\vcentcolon=\;\prod_{j=1}^{d}\Bigl[\;{\mathcal{H}}_{\beta_{\ell j}^{\star}}\!\bigl(q;\bar{p}_{\ell j},\,m_{\ell}\bigr)\;\Bigr]^{w_{\ell j}}.$
Define the edge–aggregated source weights
$M_{i\ell}({\bm{P}})\;\vcentcolon=\;\sum_{j=1}^{n}{\bm{W}}_{ij}\,{\bm{P}}_{i\ell}\,\bigl(1-{\bm{P}}_{j\ell}\bigr),$
so that the total contribution of cluster $\ell$ is:
$U_{\ell}({\bm{P}})\vcentcolon=\sum_{i=1}^{n}M_{i\ell}({\bm{P}})\,\Phi_{\ell}\!\bigl(s_{i}\bigr),$ (26)
and $U({\bm{P}})\vcentcolon=\sum_{\ell=1}^{k}U_{\ell}({\bm{P}})$. By construction (linearity of expectation and Theorem 3.10), $I_{\mathrm{true}}\!\leq U$.
Within each bin, replacing $\{{\bm{P}}_{u\ell}\}_{u\in S_{\ell j}}$ by their mean $\bar{p}_{\ell j}$ induces an AM–GM gap controlled by Propositions 3.6 and 3.8. A conservative, separable upper bound for cluster $\ell$ is:
$\Gamma_{\ell}({\bm{P}})\vcentcolon=\sum_{i=1}^{n}M_{i\ell}({\bm{P}})\,\Biggl[\;\sum_{j=1}^{d}w_{\ell j}\,{\mathbb{A}}(\beta_{\ell j}^{\star},{\bm{p}}_{\ell j},m_{\ell})\Biggr],$ (27)
and $\mathbb{A}$} is the zero–aware coefficient fromProposition 3.8. Summing over clusters, $\Gamma({\bm{P}})\vcentcolon=\sum_{\ell=1}^{k}\Gamma_{\ell}({\bm{P}})$ satisfies
$0\;\leq\;U({\bm{P}})-I_{\mathrm{true}}({\bm{P}})\;\leq\;\Gamma({\bm{P}}).$
We minimize a penalized majorizer of the expected GraphCut:
$\boxed{\min_{\;{\bm{P}}={\bm{P}}({\bm{z}})}{\mathbb{J}}_{\rho}({\bm{P}})\vcentcolon=U({\bm{P}})\;+\;\rho\,\Gamma({\bm{P}}),\quad\rho\geq 0,\quad}$ (28)
where $\bm{z}$} can be the parameterization logits (via Softmax). Since $I_{\mathrm{true}}\leq U$ and $\Gamma\geq 0$, we retain $I_{\mathrm{true}}\leq{\mathbb{J}}_{\rho}$ for all $\rho\geq 0$ while explicitly shrinking the AM–GM gap.
We detail in  Appendix B the forward-backward derivation and implementation for our final objective.
3.5 Optimization properties
We conclude here with several properties that can be useful for various extensions and gradient-based approaches. Recall the binwise envelope ${\mathcal{H}}_{b}(q;\bar{p},m)$ and the per-cluster factor $\Phi_{\ell}(q)=\prod_{j=1}^{d}\bigl[{\mathcal{H}}_{b_{\ell j}^{\star}}(q;\bar{p}_{\ell j},m_{\ell})\bigr]^{w_{\ell j}}$, with $w_{\ell j}=m_{\ell j}/m_{\ell}$, $q=s_{i}$, $c_{\ell j}\vcentcolon=\tfrac{q}{b_{\ell j}^{\star}}+1$.
Lemma 3.11 (Bounds, monotonicity, curvature).
Fix $m\in\mathbb{N}$, $b>0$, $q>0$, $c\vcentcolon=q/b+1>1$, and $z\in[0,1]$.
  1. 1.
    Second derivative of $\mathcal{H}$} is bounded:
    $0\leq\partial_{zz}{\mathcal{H}}_{b}(q;z,m)\leq\frac{1}{q}\cdot\frac{2m(m-1)}{c(c+1)}.$
  2. 2.
    Log-derivative bound: For all $z\in[0,1]$,
    $\left|\partial_{z}\log{\mathcal{H}}_{b}(q;z,m)\right|\;=\;\frac{|\partial_{z}{\mathcal{H}}_{b}|}{{\mathcal{H}}_{b}}\;\leq\;\frac{m(c+m-1)}{c(c-1)}.$
Proposition 3.12 (Properties $U_{\ell}$).
Fix $\ell$. With $\bar{{\bm{p}}}_{\ell}\!=\!(\bar{p}_{\ell 1},\dots,\bar{p}_{\ell d})$ held as variables and all other quantities fixed:
  1. 1.
    Monotone decrease in each bin mean: $\partial_{\bar{p}_{\ell j}}U_{\ell}({\bm{P}})\leq 0$ for every $j$.
  2. 2.
    Separate convexity: As a univariate function of $\bar{p}_{\ell j}$ (others fixed), $U_{\ell}$ is convex on $[0,1]$.
  3. 3.
    Lipschitz in bin means: For each $j$,
    $\left|\partial_{\bar{p}_{\ell j}}U_{\ell}({\bm{P}})\right|\;\leq\;\operatorname{Cut}_{\ell}({\bm{P}})\;w_{\ell j}\;\frac{m_{\ell}(c_{\ell j}+m_{\ell}-1)}{c_{\ell j}(c_{\ell j}-1)},$
    where $\operatorname{Cut}_{\ell}({\bm{P}})=\sum_{i,v}{\bm{W}}_{iv}\,{\bm{P}}_{i\ell}(1-{\bm{P}}_{v\ell})$ and $c_{\ell j}=\tfrac{q}{b_{\ell j}^{\star}}+1$ with $q=s_{i}$ bounded below by $q_{\min}>0$.
  4. 4.
    Coordinate Lipschitz in entries of $\bm{P}$}: If $u\in S_{\ell j}$ then $\partial\bar{p}_{\ell j}/\partial{\bm{P}}_{u\ell}=1/m_{\ell j}$. Hence:
    $\left|\partial_{{\bm{P}}_{u\ell}}U_{\ell}({\bm{P}})\right|$ $\;\leq\;\underbrace{\left(\sum_{v}{\bm{W}}_{uv}(1-{\bm{P}}_{v\ell})\right)}_{\leq d_{u}}\,\Phi_{\ell}(s_{u})$
    $\;+\;\operatorname{Cut}_{\ell}({\bm{P}})\;\frac{w_{\ell j}}{m_{\ell j}}\;\frac{m_{\ell}(c_{\ell j}+m_{\ell}-1)}{c_{\ell j}(c_{\ell j}-1)}.$
    Since $\Phi_{\ell}(s_{u})\leq\frac{1}{q_{\min}}$ by Lemma 3.11(1), the RHS is explicit.
Lemma 3.13 (Convexity and Lipschitzness of the variance term).
For each bin $S_{\ell j}$:
$\mathrm{Var}_{\ell j}(p)\;=\;\frac{1}{m_{\ell j}}\sum_{u\in S_{\ell j}}\bigl({\bm{P}}_{u\ell}-\bar{p}_{\ell j}\bigr)^{2}$
is convex in the vector $({\bm{P}}_{u\ell})_{u\in S_{\ell j}}$, with $\nabla_{{\bm{P}}_{u\ell}}\mathrm{Var}_{\ell j}(p)=\tfrac{2}{m_{\ell j}}({\bm{P}}_{u\ell}-\bar{p}_{\ell j})$, and $\|\nabla\mathrm{Var}_{\ell j}\|_{\infty}\leq\tfrac{2}{m_{\ell j}}$.
$\Gamma_{\ell}({\bm{P}})$ is a product of a globally Lipschitz function and a convex nonnegative function $\sum_{j}w_{\ell j}{\mathbb{A}}_{\ell j}\mathrm{Var}_{\ell j}(p)$. Although the product need not be convex, for any $\eta>0$, Young’s inequality gives the convex upper bound
$\Gamma_{\ell}({\bm{P}})\leq\!\sum_{j}\frac{m_{\ell}w_{\ell j}}{4}{\mathbb{A}}_{\ell j}\Bigl(\eta\,\operatorname{Cut}_{\ell}({\bm{P}})^{2}+\tfrac{1}{\eta}\mathrm{Var}_{\ell j}(p)^{2}\Bigr).$
Thus the penalized objective admits a convex majorizer that can be minimized each iteration (MM), guaranteeing monotone descent of the true ${\mathbb{J}}_{\rho}$.
Time complexity.
With minibatches of size $B$, we first construct the batch adjacency ${\bm{W}}_{\text{batch}}\in\mathbb{R}^{B\times B}_{+}$. For dense similarities this costs $O(B^{2})$ time (and $O(B^{2})$ memory); in sparse/kkNN settings replace $B^{2}$ by $\mathrm{nnz}({\bm{W}}_{\text{batch}})$. We then precompute the envelope terms ${\mathcal{H}}_{\beta^{\star}_{j},\ell}$ for every (bin $j$, cluster $\ell$). A straightforward implementation performs O(dkm)O(d\,k\,m) work, where $d$ is the number of bins, $k$ the number of clusters, and $m$ the polynomial degree in the $_{2}F_{1}(-m,\cdot\,;\cdot\,;z)$ evaluation. Because these computations factor across $(j,\ell)$, they are embarrassingly parallel; with $(d\times k)$ workers the wall-clock reduces to $O(m)$ (see Appendix B). Thus, each batch step would practically take $O(B^{2}k+m)$
4 Related Work
The reader familiar with Spectral Clustering wonders why the graph Laplacian does not appear in our soft assignment formulation. Let ${\bm{f}}\in\{0,1\}^{n}$ denote the indicator of a cluster $\mathbb{C}$}, and let ${\bm{p}}\in[0,1]^{n}$ with $p_{i}\vcentcolon=\Pr[{\bm{f}}_{i}=1]$. Then:
$\mathbb{E}\left[{\bm{f}}^{\top}{\bm{D}}{\bm{f}}\right]\;=\;\sum_{i=1}^{n}d_{i}\,p_{i};\quad{\bm{p}}^{\top}{\bm{D}}{\bm{p}}\;=\;\sum_{i=1}^{n}d_{i}\,p_{i}^{\,2}\,.$ (29)
Hence ${\bm{p}}^{\top}{\bm{D}}{\bm{p}}\leq\mathbb{E}\left[{\bm{f}}^{\top}{\bm{D}}{\bm{f}}\right]$ with equality if and only if $p_{i}\in\{0,1\}$ for all $i$. Consequently, the Laplacian quadratic ${\bm{p}}^{\top}{\bm{L}}{\bm{p}}$ coincides with the expected cut form $\mathbb{E}\left[{\bm{f}}^{\top}{\bm{L}}{\bm{f}}\right]$ only in the hard (binary) case.
Indeed, for ${\bm{f}}_{i}\in\{0,1\}$, we have ${\bm{f}}_{i}^{2}={\bm{f}}_{i}$, and thus $\mathbb{E}\left[{\bm{f}}^{\top}{\bm{D}}{\bm{f}}\right]=\sum_{i}d_{i}\,\mathbb{E}\left[{\bm{f}}_{i}]=\sum_{i}d_{i}\,p_{i}\right]$ (no independence assumptions are needed). On the other hand, for any ${\bm{p}}\in[0,1]^{n}$, ${\bm{p}}^{\top}{\bm{D}}{\bm{p}}=\sum_{i}d_{i}\,p_{i}^{2}\leq\sum_{i}d_{i}\,p_{i}=\mathbb{E}\left[{\bm{f}}^{\top}{\bm{D}}{\bm{f}}\right]$ because $p_{i}^{2}\leq p_{i}$ on $[0,1]$, with equality iff $p_{i}\in\{0,1\}$ for all $i$. For the adjacency term, $\mathbb{E}\left[{\bm{f}}^{\top}{\bm{W}}{\bm{f}}\right]=\sum_{i,j}{\bm{W}}_{ij}\,\mathbb{E}\left[{\bm{f}}_{i}{\bm{f}}_{j}\right]$, which equals ${\bm{p}}^{\top}{\bm{W}}{\bm{p}}$ only under the additional assumption $\mathbb{E}\left[{\bm{f}}_{i}{\bm{f}}_{j}\right]=p_{i}p_{j}$ (e.g., independence). Even under independence, the degree-term mismatch persists unless $\bm{p}$} is binary, so ${\bm{p}}^{\top}{\bm{L}}{\bm{p}}=\mathbb{E}\left[{\bm{f}}^{\top}{\bm{L}}{\bm{f}}\right]$ holds only in the hard case.
In the hard case with ${\bm{f}}={\bm{1}}_{{\mathbb{C}}}$, one has ${\bm{f}}^{\top}{\bm{D}}{\bm{f}}=\sum_{i}d_{i}{\bm{f}}_{i}=\operatorname*{vol}\left({\mathbb{C}}\right)$ and ${\bm{f}}^{\top}{\bm{L}}{\bm{f}}$ equals the standard cut expression. In the soft case, the probabilistic objective naturally produces the linear degree contribution $\sum_{i}d_{i}p_{i}$ (expected volume) rather than the quadratic $\sum_{i}d_{i}p_{i}^{\,2}$ required by the Laplacian quadratic. Therefore, replacing the expected degree term by ${\bm{p}}^{\top}{\bm{D}}{\bm{p}}$ in a simplex relaxation agrees on the binary vertices but mis-specifies the interior and can change the optimizer. Our formulation keeps the expected (linear) degree term and does not assert a Laplacian quadratic unless assignments are binary.
4.1 A cross-entropy upper bound and the contrastive link
For $p\in(0,1]$, $1-p\leq-\log p$. Hence, for any edge $(i,j)$ and cluster $\ell$:
${\bm{P}}_{i\ell}(1-{\bm{P}}_{j\ell})+{\bm{P}}_{j\ell}(1-{\bm{P}}_{i\ell})\leq{\bm{P}}_{i\ell}\!\big(-\log{\bm{P}}_{j\ell}\big)+{\bm{P}}_{j\ell}\!\big(-\log{\bm{P}}_{i\ell}\big).$
Summing over $\ell$ gives a bound by the symmetric cross-entropy $\text{CE}({\bm{P}}_{i}\!\|{\bm{P}}_{j})+\text{CE}({\bm{P}}_{j}\!\|{\bm{P}}_{i})$; summing further with weights ${\bm{W}}_{ij}$ upper-bounds the expected-cut terms in equation 3 (and the numerators of equation 1).
Let ${\bm{P}}_{i\ell}=\mathrm{softmax}_{\ell}(z_{i\ell}/\tau)$ with logits $z_{i\ell}$ and temperature $\tau>0$. Then $\text{CE}({\bm{P}}_{i}\!\|{\bm{P}}_{j})=H({\bm{P}}_{i})+D_{\mathrm{KL}}({\bm{P}}_{i}\!\|{\bm{P}}_{j})$, so minimizing the symmetric surrogate encourages agreement between ${\bm{P}}_{i},{\bm{P}}_{j}$ and lower entropy (sharper assignments). If the index $\ell$ ranges over batch instances (keys), then $\text{CE}({\bm{P}}_{i}\!\|{\bm{P}}_{j})=-\log{\bm{P}}_{j,i}$ recovers the InfoNCE term; using both directions gives the standard symmetric SimCLR/CLIP loss. If $\ell$ indexes prototypes, the same surrogate enforces code consistency (prototypical contrast). General ${\bm{W}}_{ij}\!\geq 0$ yields a graph-weighted variant that selects positives via the affinity graph.
Smaller $\tau$ sharpens $\bm{P}$} and narrows the bound; larger $\tau$ smooths codes but loosens it. The bound is tightest when neighbors share peaky assignments; it is looser for diffuse or mismatched neighborhoods.
4.2 SimCLR as a special case of our framework
Let $x_{i}$ be an image and $(x_{i}^{(1)},x_{i}^{(2)})$ two augmentations. SimCLR computes embeddings $z_{i}^{(1)}=f_{\theta}(x_{i}^{(1)})$ and $z_{i}^{(2)}=f_{\theta}(x_{i}^{(2)})$, uses a temperatured cosine similarity $\kappa(u,v)=\exp(\langle u,v\rangle/\tau)$, and the InfoNCE loss where $(i,i^{+})$ is the positive pair and the rest are negatives. Build a view graph on all augmented samples with
${\bm{W}}_{ab}\;\vcentcolon=\;\kappa\!\bigl(z_{a},z_{b}\bigr),$
and take RatioCut ($s_{i}\equiv 1$) or Ncut ($s_{i}=d_{i}$). Set $k$ to the number of latent classes; in the extreme “instance discrimination” formulation ($k$ equals the batch size), each instance forms its own cluster.
There is a single representation space, so we use a single bin (d=1d{=}1), $\beta^{\star}$ shared across vertices (Section 3.3). The per-cluster envelope is:
$U_{\ell}({\bm{P}})\;=\;\sum_{i=1}^{n}\Bigl[\sum_{v}{\bm{W}}_{iv}\,{\bm{P}}_{i\ell}\,(1-{\bm{P}}_{v\ell})\Bigr]\,{\mathcal{H}}\!\bigl(s_{i};\,\bar{p}_{\ell},\,m_{\ell}\bigr),$
with $\bar{p}_{\ell}=m_{\ell}^{-1}\sum_{u\in{\mathbb{C}}_{\ell}}{\bm{P}}_{u\ell}$. Since $\mathcal{H}$} is decreasing in $\bar{p}$ (Lemma 3.4), increasing same-class agreement (larger $\bar{p}_{\ell}$) monotonically decreases $U_{\ell}$.
As SimCLR improves alignment (increases $\kappa$ for positives) and uniformity (decreases $\kappa$ for negatives), the expected cross-edge mass $\sum_{(i,j):y_{i}\neq y_{j}}{\bm{W}}_{ij}$ decreases in the view graph built from $\kappa$. This is the standard interpretive link for InfoNCE: positives up/negatives down implies fewer large-weight cross edges in the similarity graph.
Fix the similarity $\kappa$ and graph $\bm{W}$} built from the same embeddings as SimCLR in a given iteration. For $d=1$, RatioCut or Ncut, and any $\rho\geq 0$, minimizers of our gap-aware objective ${\mathbb{J}}_{\rho}({\bm{P}})=U({\bm{P}})+\rho\,\Gamma_{\text{simple}}({\bm{P}})$ select labelings that (for fixed embeddings) improve alignment and uniformity. Consequently, under the encoder updates that drive InfoNCE down, ${\mathbb{J}}_{\rho}$ decreases monotonically as well; in the instance-discrimination limit ($k$ equals batch size), the minimizers coincide with those of SimCLR up to the order-preserving reparameterization ${\bm{P}}=\mathrm{softmax}({\bm{z}})$.
For fixed embeddings, reducing cross-edge mass reduces the cut term $M_{i\ell}=\sum_{v}{\bm{W}}_{iv}{\bm{P}}_{i\ell}(1-{\bm{P}}_{v\ell})$, hence reduces $U_{\ell}$ because $\mathcal{H}$} is positive and independent of $(i,v)$. InfoNCE steps that increase positive $\kappa$ and decrease negative $\kappa$ strictly decrease $\sum_{(i,v):y_{i}\neq y_{v}}{\bm{W}}_{iv}$, so $\sum_{\ell}U_{\ell}$ decreases. The zero-aware variance penalty $\Gamma_{\text{simple}}$ is nonnegative and discourages collapse, analogous to SimCLR’s temperature parameter that pushes the probabilities to the binary case as the temperature decreases.
Practical recipe.
Build $\bm{W}$} with the same $\kappa$ and temperature $\tau$ as SimCLR; use $d=1$; and then optimize $\min_{{\bm{P}}}\log U({\bm{P}})$ (the $\log$ is order-preserving, makes the product-to-sum trivial here). Decay the temperature to a low target as a heuristic too minimize $\rho\,\Gamma({\bm{P}})$.
4.3 CLIP as a probabilistic graph cut
[…]
${\bm{W}}_{ij}^{xt}\;\vcentcolon=\;\kappa\!\bigl(z_{i},u_{j}\bigr),\qquad{\bm{W}}=\begin{pmatrix}0&{\bm{W}}^{xt}\\ ({\bm{W}}^{xt})^{\top}&0\end{pmatrix}$.
Use RatioCut ($s\equiv 1$) or Ncut ($s_{i}=d_{i}$). We define $k$ text-conditioned clusters: each text node $t_{j}$ indexes a cluster $\ell=j$, and the model learns soft assignments ${\bm{P}}_{i\ell}$ of images to text clusters (and symmetrically, soft assignments of texts to image clusters).
There are two representation spaces (image, text). Use $d=2$ bins with representatives $\beta_{x}^{\star}$ and $\beta_{t}^{\star}$ chosen as in Sec. 3.3 to preserve the bound direction. The per-cluster envelope becomes the product of the two single-space envelopes (Hölder with $p_{j}=m/m_{j}$).
As CLIP increases image–text alignment for matched pairs and decreases similarity for unmatched pairs, the cross-partition mass $\sum_{(i,j):\text{mismatch}}{\bm{W}}_{ij}^{xt}$ decreases in the bipartite graph built from $\kappa$.
[…]
Practical recipe.
Build the bipartite $\bm{W}$} with the same temperature and cosine similarity as CLIP; use $d=2$ bins (image/text) with left-endpoint representatives to keep the bound valid; then train with $\min_{{\bm{P}}}\log U({\bm{P}})+\rho\,\Gamma_{\text{simple}}({\bm{P}})$; the log yields a sum over modalities (matching CLIP’s two-headed loss) while keeping minimizers unchanged.
5 Conclusion
We introduced a unified, probabilistic framework for differentiable graph partitioning that yields tight, analytic upper bounds on expected discrete cuts, including Normalized Cut, with closed-form forward and backward via $_{2}F_{1}$ primitives. By moving beyond RatioCut-only relaxations, we establish a numerically stable surrogate that is provably monotone in the sufficient statistics (bin means), separately convex in each bin, and endowed with explicit Lipschitz/curvature constants for principled first-order optimization. On top of the envelope, we quantify and control the AM–GM/Jensen gap through two-sided bounds and a zero-aware dispersion penalty, prove minibatch concentration for the hypergeometric estimator, and handle heterogeneous degrees via a Hölder-product binning scheme that remains near-tight under a colinearity criterion. These ingredients together deliver a practical, theory-backed objective for scalable, end-to-end learning with graph cuts and connect directly to widely used SSL pipelines (e.g., SimCLR/CLIP) when the graph is built from current embeddings.
Limitations and scope.
Our analysis assumes independent Bernoulli assignments within clusters; modeling dependencies (e.g., submodular or MRF couplings) remains open. The Hölder envelope is tight under near-colinearity across bins; extremely heavy-tailed degree distributions may require finer, adaptive binning. Finally, while $_{2}F_{1}$ evaluation is stable in our setting (a=ma=-m), very large $m$ benefits from numerically guarded recurrences or compensated summation.
Future directions.
Three directions follow naturally: Adaptive, data-driven binning that optimizes a colinearity proxy online; Dependent relaxations (e.g., exchangeable or DPP priors) to capture richer assignment structure while retaining tractable envelopes; and Task-coupled graphs where similarities are learned jointly with $\bm{P}$}, unifying graph construction and cut optimization under a single objective. We expect these extensions to broaden the reach of differentiable cut objectives across clustering, metric learning, and multimodal pretraining.
Appendix A Proofs
A.1 Generalized Poisson-Binomial
Lemma A.1 (PGF of a weighted Bernoulli sum).
Let $r_{i}\sim\mathrm{Bernoulli}(\alpha_{i})$ be independent and define $X\vcentcolon=\sum_{i=1}^{m}\beta_{i}r_{i}$ with $\beta_{i}\in\mathbb{Z}_{\geq 0}$. Then the probability generating function $G_{X}(t)\vcentcolon=\mathbb{E}[t^{X}]$ (for $|t|\leq 1$) is
$G_{X}(t)\;=\;\prod_{i=1}^{m}\bigl((1-\alpha_{i})+\alpha_{i}\,t^{\beta_{i}}\bigr).$
Proof.
Since $X=\sum_{i}\beta_{i}r_{i}$ and $r_{i}\in\{0,1\}$, $t^{X}=\prod_{i=1}^{m}t^{\beta_{i}r_{i}}.$ By independence,
$G_{X}(t)\;=\;\mathbb{E}\!\left[\prod_{i=1}^{m}t^{\beta_{i}r_{i}}\right]=\prod_{i=1}^{m}\mathbb{E}\!\left[t^{\beta_{i}r_{i}}\right].$
For each $i$, $\mathbb{E}[t^{\beta_{i}r_{i}}]=(1-\alpha_{i})t^{0}+\alpha_{i}t^{\beta_{i}}=(1-\alpha_{i})+\alpha_{i}t^{\beta_{i}}$. Multiplying the factors yields the claim. ∎
A.2 Integral Representation: Proof of Lemma 3.2
See 3.2
Proof.
The proof uses the integral representation of the reciprocal. For any $X>0$, we have $\frac{1}{X}=\int_{0}^{1}t^{X-1}\,dt$. Applying this with $X=q+{\textnormal{x}}$ (which is a.s. positive since $q>0$ and ${\textnormal{x}}\geq 0$),
$\frac{1}{q+{\textnormal{x}}}=\int_{0}^{1}t^{\,q+{\textnormal{x}}-1}\,dt.$
Taking expectations and using Tonelli’s theorem (the integrand $t^{\,q+{\textnormal{x}}-1}$ is nonnegative on $[0,1]$),
$\mathbb{E}\!\left[\frac{1}{q+{\textnormal{x}}}\right]$ $=\mathbb{E}\!\left[\int_{0}^{1}t^{\,q+{\textnormal{x}}-1}\,dt\right]=\int_{0}^{1}\mathbb{E}\!\left[t^{\,q+{\textnormal{x}}-1}\right]dt$
$=\int_{0}^{1}\mathbb{E}\!\left[t^{\,q-1}\,t^{{\textnormal{x}}}\right]dt=\int_{0}^{1}t^{\,q-1}\,\mathbb{E}\!\left[t^{{\textnormal{x}}}\right]dt.$
Here $\mathbb{E}[t^{{\textnormal{x}}}]$ is the probability generating function (PGF) of x, denoted $G_{{\textnormal{x}}}(t)$. Substituting the PGF from Equation 6 gives
$\mathbb{E}\!\left[\frac{1}{q+{\textnormal{x}}}\right]=\int_{0}^{1}t^{\,q-1}G_{{\textnormal{x}}}(t)\,dt=\int_{0}^{1}t^{\,q-1}\!\left[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta_{i}}\bigr)\right]dt.$
A.3 Hypergeometric Bound: Proof of Theorem 3.5
See 3.5
Proof.
Assume $\beta_{i}\equiv\beta>0$ and $q>0$. Recall the definition of ${\mathcal{I}}(q,{\bm{\alpha}},\beta)$:
${\mathcal{I}}(q,{\bm{\alpha}},\beta)\vcentcolon=\int_{0}^{1}\!\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\Biggr]t^{q-1}\,dt.$
For fixed $t\in[0,1]$, the map $\alpha\mapsto\log\!\bigl(1-\alpha+\alpha t^{\beta}\bigr)$ is concave (log of a positive affine function), hence by Jensen:
$\sum_{i=1}^{m}\log\!\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\;\leq\;m\log\!\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr),\quad\bar{\alpha}\vcentcolon=\tfrac{1}{m}\sum_{i=1}^{m}\alpha_{i}.$
Exponentiating and integrating gives:
$I\;\leq\;\int_{0}^{1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}t^{q-1}\,dt\;=\;B.$
Evaluate $B$ via $u=t^{\beta}$ (so $dt=\tfrac{1}{\beta}u^{\tfrac{1}{\beta}-1}du$):
$B=\frac{1}{\beta}\int_{0}^{1}(1-\bar{\alpha}+\bar{\alpha}u)^{m}\,u^{\tfrac{q}{\beta}-1}\,du=\frac{1}{\beta}\int_{0}^{1}(1-\bar{\alpha}v)^{m}(1-v)^{\tfrac{q}{\beta}-1}\,dv,$
with v=1uv=1-u. By Euler’s integral for $_{2}F_{1}$ with $(a,b,c,z)=(-m,1,\tfrac{q}{\beta}+1,\bar{\alpha})$ (valid since $c>b>0$),
$B=\frac{1}{\beta}\cdot\frac{\Gamma(1)\Gamma(\tfrac{q}{\beta})}{\Gamma(\tfrac{q}{\beta}+1)}\,{}_{2}F_{1}\!\Bigl(-m,1;\tfrac{q}{\beta}+1;\bar{\alpha}\Bigr)=\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;\tfrac{q}{\beta}+1;\bar{\alpha}\Bigr).$
Therefore:
$\boxed{\ I\;\leq\;\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;\tfrac{q}{\beta}+1;\bar{\alpha}\Bigr)\ }.$
A.4 AM-GM Gap: Proof of Proposition 3.6
See 3.6
Proof.
Fix $\tau\in[0,1]$ and set $c\vcentcolon=1-\tau$. Define $f_{\tau}(\alpha)\vcentcolon=\log(1-\alpha+\alpha\tau)=\log(1-c\,\alpha)$, a concave function on $[0,1]$ with:
$f_{\tau}^{\prime\prime}(\alpha)=-\,\frac{c^{2}}{(1-c\,\alpha)^{2}}\ \in\ \Bigl[-\frac{c^{2}}{\tau^{2}},\,-\,c^{2}\Bigr].$
Thus $f_{\tau}$ is $L_{\tau}$-smooth with $L_{\tau}=c^{2}/\tau^{2}$ and $\gamma_{\tau}$-strongly convex with $\gamma_{\tau}=c^{2}$ on $[0,1]$. By the standard Jensen two-sided bound for twice-differentiable concave functions:
$\frac{\gamma_{\tau}}{2}\,\mathrm{Var}({\bm{\alpha}})\;\leq\;f_{\tau}(\bar{\alpha})-\frac{1}{m}\sum_{i=1}^{m}f_{\tau}(\alpha_{i})\;\leq\;\frac{L_{\tau}}{2}\,\mathrm{Var}({\bm{\alpha}}).$ (30)
Exponentiating equation 30 yields the pointwise multiplicative AM–GM control (for any fixed $\tau\in[0,1]$):
$\exp\!\Bigl(\tfrac{m}{2}\gamma_{\tau}\,\mathrm{Var}({\bm{\alpha}})\Bigr)\;\leq\;\frac{\bigl(1-\bar{\alpha}+\bar{\alpha}\,\tau\bigr)^{m}}{\prod_{i=1}^{m}(1-\alpha_{i}+\alpha_{i}\tau)}\;\leq\;\exp\!\Bigl(\tfrac{m}{2}L_{\tau}\,\mathrm{Var}({\bm{\alpha}})\Bigr),$ (31)
with equality iff $\mathrm{Var}({\bm{\alpha}})=0$ (or $\tau\in\{0,1\}$). Equivalently, the additive gap satisfies;
$\bigl(1-\bar{\alpha}+\bar{\alpha}\,\tau\bigr)^{m}\!\Bigl(1-e^{-\frac{m}{2}L_{\tau}\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\;\geq\;\bigl(1-\bar{\alpha}+\bar{\alpha}\,\tau\bigr)^{m}-\!\!\prod_{i=1}^{m}(1-\alpha_{i}+\alpha_{i}\tau)\;\geq\;\bigl(1-\bar{\alpha}+\bar{\alpha}\,\tau\bigr)^{m}\!\Bigl(1-e^{-\frac{m}{2}\gamma_{\tau}\,\mathrm{Var}({\bm{\alpha}})}\Bigr).$ (32)
Now set $\tau=t^{\beta}$ (the theorem’s common exponent) and multiply equation 32 by $t^{q-1}$, then integrate $t\in[0,1]$. Using:
$I(q;{\bm{\alpha}},\beta)\ \vcentcolon=\ \int_{0}^{1}\!\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\Biggr]t^{q-1}\,dt,\quad B_{\mathrm{AMGM}}(q)\ \vcentcolon=\ \int_{0}^{1}\!\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}t^{q-1}\,dt,$
we obtain the deterministic two-sided bound:
$\underline{\Delta}(q)\;\leq\;B_{\mathrm{AMGM}}(q)\;-\;I(q;{\bm{\alpha}},\beta)\;\leq\;\overline{\Delta}(q),$ (33)
$\underline{\Delta}(q)\vcentcolon=\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}\Bigl(1-e^{-\frac{m}{2}\,\gamma(t)\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
$\overline{\Delta}(q)\vcentcolon=\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}\Bigl(1-e^{-\frac{m}{2}\,L(t)\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
where $\gamma(t)\vcentcolon=(1-t^{\beta})^{2}$ and $L(t)\vcentcolon=\gamma(t)/t^{2\beta}$, and $B_{\mathrm{AMGM}}(q)\vcentcolon={\mathcal{H}}_{\beta}(q,\bar{\alpha},\beta)$.
Using $1-e^{-x}\leq x$ gives the simple upper bound:
$0\;\leq\;B_{\mathrm{AMGM}}(q)-I(q;{\bm{\alpha}},\beta)\;\leq\;\frac{m}{2}\,\mathrm{Var}({\bm{\alpha}})\!\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}L(t)\,dt,$ (34)
and $1-e^{-x}\geq\tfrac{x}{1+x}$ yields a corresponding explicit lower bound. The bounds in equation 33 are tight iff $\mathrm{Var}({\bm{\alpha}})=0$, in which case $B_{\mathrm{AMGM}}(q)=I(q;{\bm{\alpha}},\beta)$. ∎
A.5 Proofs for 0-aware gap
See 3.8
Proof.
Fix $t\in[0,1]$ and write $\tau\vcentcolon=t^{\beta}\in[0,1]$. Let $f_{\tau}(\alpha)\vcentcolon=\log(1-\alpha+\alpha\tau)$, which is concave on $[0,1]$ with
$f_{\tau}^{\prime\prime}(\alpha)=\frac{(1-\tau)^{2}}{(1-(1-\tau)\alpha)^{2}}\in\Bigl[(1-\tau)^{2},\,(1-\tau)^{2}/\tau^{2}\Bigr].$
Set the zero-aware weights $\lambda_{i}\vcentcolon=\omega_{0}(\alpha_{i})/\Omega$ and denote the $\omega_{0}$–weighted mean and variance by $\bar{\alpha}^{\omega_{0}}=\sum_{i}\lambda_{i}\alpha_{i}$ and $\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})=\sum_{i}\lambda_{i}(\alpha_{i}-\bar{\alpha}^{\omega_{0}})^{2}$ (with the usual convention $\mathrm{Var}^{\omega_{0}}=0$ if $\Omega=0$).
By weighted Jensen for the concave $f_{\tau}$:
$\sum_{i=1}^{m}\lambda_{i}f_{\tau}(\alpha_{i})\ \leq\ f_{\tau}(\bar{\alpha}^{\omega_{0}}).$
The standard second-order (weighted) Jensen gap bound gives:
$f_{\tau}(\bar{\alpha}^{\omega_{0}})-\sum_{i=1}^{m}\lambda_{i}f_{\tau}(\alpha_{i})\ \leq\ \frac{L_{\tau}}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}}),\qquad L_{\tau}\vcentcolon=\frac{(1-\tau)^{2}}{\tau^{2}}.$
Multiplying by $\Omega$ and exponentiating yields the pointwise zero-aware AM–GM control:
[…]
where $Y(t)\!=\!\prod_{i}(1-\alpha_{i}+\alpha_{i}t^{\beta})$ and $\bar{Y}(t)\!=\!(1-\bar{\alpha}+\bar{\alpha}t^{\beta})^{m}$ (note: we keep the envelope centered at the plain mean $\bar{\alpha}$ as in Theorem 3.5).
Using $1-e^{-x}\leq x$ and integrating against $t^{q-1}$ gives:
$0\ \leq\ B_{\mathrm{AMGM}}(q)-{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \frac{m}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\,\int_{0}^{1}t^{q-1}\bar{Y}(t)\,\frac{(1-t^{\beta})^{2}}{t^{2\beta}}\,dt.$
By differentiating under the integral sign and the binomial identity $\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}t^{r\beta}=(1-t^{\beta})^{2}$, one obtains the identity (derived in the main text):
$\int_{0}^{1}t^{q-1}\bar{Y}(t)\,\frac{(1-t^{\beta})^{2}}{t^{2\beta}}\,dt\;=\;\frac{\partial}{\partial\bar{\alpha}}\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\,{\mathcal{H}}_{\beta}\!\bigl(q+r\beta;\bar{\alpha},m\bigr)\;=\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr),$
valid for $q>2\beta$ by Euler’s integral (and for all $q>0$ by analytic continuation). Combining the two displays yields
$B_{\mathrm{AMGM}}(q)-{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \frac{m}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr)\ \vcentcolon=\ {\mathbb{A}}(q,{\bm{\alpha}},m),$
which is exactly equation 16. The bound is zero-aware since $\omega_{0}(0)=0$ removes inactive coordinates from $\mathrm{Var}^{\omega_{0}}$; it is tight when $\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})=0$ (i.e., the $\omega_{0}$–weighted dispersion vanishes), in which case $\bar{Y}(t)\equiv Y(t)$ and equality holds. ∎
A.6 Hölder product bound for heterogeneous exponents
Let ${\bm{\beta}}=(\beta_{i})_{i=1}^{m}$ and take $d$ distinct values $\{b_{1},\dots,b_{d}\}\subset(0,\infty)$, and partition indices by $S_{j}\vcentcolon=\{i:\beta_{i}=b_{j}\}$ with sizes $m_{j}\vcentcolon=|S_{j}|$ and $\sum_{j=1}^{d}m_{j}=m$. Define the group means:
$\bar{\alpha}_{j}\;\vcentcolon=\;\frac{1}{m_{j}}\sum_{i\in S_{j}}\alpha_{i}\in[0,1].$
Recall the objective integral:
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \vcentcolon=\ \int_{0}^{1}\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta_{i}}\bigr)\Biggr]\,t^{q-1}\,dt,\qquad q>0.$
Lemma A.2 (Hölder–binned envelope).
With the notation above,
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\;\leq\;\prod_{j=1}^{d}\Bigl[\ {\mathcal{H}}_{b_{j}}\bigl(q;\bar{\alpha}_{j},m\bigr)\ \Bigr]^{\,m_{j}/m},$
[…]
Proof.
For each group $S_{j}$ (fixed $t\in[0,1]$):
$P_{j}(t)\ \vcentcolon=\ \prod_{i\in S_{j}}\bigl(1-\alpha_{i}+\alpha_{i}t^{b_{j}}\bigr)\ \leq\ \bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\quad\text{by AM--GM.}$
Multiplying over $j$ gives:
$\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta_{i}}\bigr)\ \leq\ \prod_{j=1}^{d}\bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}.$
Hence:
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \int_{0}^{1}\prod_{j=1}^{d}\bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\,t^{q-1}\,dt.$
Let $w_{j}\vcentcolon=m_{j}/m$ and split $t^{q-1}=\prod_{j=1}^{d}t^{(q-1)w_{j}}$. Set:
$g_{j}(t)\ \vcentcolon=\ \bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\,t^{(q-1)w_{j}}.$
Choose exponents $p_{j}\vcentcolon=\frac{m}{m_{j}}>1$, so that $\sum_{j=1}^{d}\frac{1}{p_{j}}=\sum_{j}\frac{m_{j}}{m}=1$. By Hölder’s inequality for products,
$\int_{0}^{1}\prod_{j=1}^{d}g_{j}(t)\,dt\ \leq\ \prod_{j=1}^{d}\Bigl(\int_{0}^{1}|g_{j}(t)|^{p_{j}}\,dt\Bigr)^{1/p_{j}}.$
But $g_{j}^{p_{j}}(t)=\bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m}\,t^{q-1}$, hence:
$\int_{0}^{1}\prod_{j=1}^{d}g_{j}(t)\,dt\ \leq\ \prod_{j=1}^{d}\Bigl(\int_{0}^{1}\bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m}\,t^{q-1}\,dt\Bigr)^{m_{j}/m}.$
For each $j$, the inner integral equals:
$\int_{0}^{1}\bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m}\,t^{q-1}\,dt\;=\;{\mathcal{H}}_{b_{j}}\bigl(q;\bar{\alpha}_{j},m\bigr)\;=\;\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;\tfrac{q}{b_{j}}+1;\bar{\alpha}_{j}\Bigr),$
by the same change-of-variables/Euler-integral used in Theorem 3.5 (valid for $q>0$, $b_{j}>0$). Combining (i)–(iii) yields the stated bound. ∎
Remarks.
If exponents are grouped into bins with ranges $b_{j}^{\leftarrow},b_{j}^{\rightarrow}]$ rather than singletons, the same proof holds after replacing $b_{j}$ by any representative $b_{j}^{\leftarrow}\leq\beta_{i}$ for $i\in S_{j}$, preserving the upper-bound direction. The bound is a weighted geometric mean of $d$ hypergeometric envelopes, with weights $m_{j}/m$, and avoids collapsing all exponents to a single conservative value.
Appendix B Forward-Backward algorithms
[…]
$\frac{d}{dz}{}_{2}F_{1}(a,b;c;z)=\frac{ab}{c}\,{}_{2}F_{1}(a+1,b+1;c+1;z),$
and the fact that for a=ma=-m (or m+1-m+1) the series truncates (finite polynomial).
B.1 Efficient computation of $_{2}F_{1}$
For a=ma=-m and $b=1$, the Gauss hypergeometric reduces to a degree-$m$ polynomial:
$_{2}F_{1}(-m,1;c;z)=\sum_{k=0}^{m}\frac{(-m)_{k}(1)_{k}}{(c)_{k}}\frac{z^{k}}{k!}=\sum_{k=0}^{m}(-1)^{k}\binom{m}{k}\frac{z^{k}}{(c)_{k}},\quad z\in[0,1].$
Although the series is finite (exact after $k=m$), in practice many tails are negligible. We use an early-exit rule at index $K$ when:
$|t_{K+1}|<\varepsilon_{\mathrm{rel}}\max\{|S_{K}|,\delta_{\mathrm{abs}}\},$
where $S_{K}$ is the current partial sum, $\varepsilon_{\mathrm{rel}}$ a relative tolerance, and $\delta_{\mathrm{abs}}$ a floor for tiny values (e.g., machine epsilon scaled). This is safe because the remaining mKm-K terms are alternating and (empirically) rapidly shrinking for $z\in[0,1]$; for reproducibility one can cap $K\leq m$.
At $z=0$ the value is $1$; near $z=1$ we rely on Horner/compensation to manage cancellation. For large $c$ (e.g., the $c=2$ relaxed envelope of §B.6), coefficients become very benign: $_{2}F_{1}(-m,1;2;z)=\sum_{k=0}^{m}(-1)^{k}\binom{m}{k}z^{k}/(k+1)!$.
Applying AM–GM within bins $S_{j}$ of equal $\beta$ (i.e., $\beta_{i}=b_{j}$) gives bin-wise polynomials after replacing $m$ by $m_{j}$ and $\bar{\alpha}$ by $\bar{\alpha}_{j}$; the product envelope’s slack is then controlled by within-bin dispersions $\{\mathrm{Var}_{j}({\bm{\alpha}})\}_{j}$.
B.2 The forward pass (Hölder envelope + derivatives)
We now give a concrete forward routine that returns the Hölder envelope $B_{\mathrm{H\ddot{o}lder}}$ and the two hypergeometric building blocks needed for the backward pass (the ratio $H^{\prime}_{j}/H_{j}$” in the gradient). The algorithm evaluates:
$B_{\mathrm{H\ddot{o}lder}}=\prod_{j=1}^{d}\Bigl[{\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)\Bigr]^{m_{j}/m},\qquad{\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)=\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;c_{j};\bar{\alpha}_{j}\Bigr),\quad c_{j}=\frac{q}{b_{j}}+1.$
We accumulate in the log domain to avoid underflow/overflow.
[…]
1:Bin indices by equal $\beta$: obtain $\{(b_{j},S_{j},m_{j},\bar{\alpha}_{j})\}_{j=1}^{d}$, with $m=\sum_{j}m_{j}$.
2:for $j=1$ to $d$ do $\triangleright$ $_{2}F_{1}(-m,1;c_{j};\bar{\alpha}_{j})$ via term ratios or Horner
3:  $c_{j}\!\leftarrow\!q/b_{j}+1$; $H_{j}\!\leftarrow\!1$; $t\!\leftarrow\!1$
4:  for $k=1$ to $m$ do
5:   $t\!\leftarrow\!t\cdot\frac{(-m+k-1)}{(c_{j}+k-1)}\cdot\bar{\alpha}_{j}$ $\triangleright$ alternates in sign
6:   $H_{j}\!\leftarrow\!H_{j}+t$ $\triangleright$ use Kahan/Neumaier compensation
7:   if $|t|<\varepsilon_{\mathrm{rel}}\max(|H_{j}|,\delta_{\mathrm{abs}})$ then break
8:  $B_{j}^{\ast}\!\leftarrow\!H_{j}/q$; $\ell_{j}\!\leftarrow\!\frac{m_{j}}{m}\log B_{j}^{\ast}$
[…]
10:for $j=1$ to $d$ do $\triangleright$ $_{2}F_{1}(-m+1,2;c_{j}+1;\bar{\alpha}_{j})$ for gradients
11:  $H^{\prime}_{j}\!\leftarrow\!1$; $t\!\leftarrow\!1$
12:  for $k=1$ to m1m-1 do
13:   $t\!\leftarrow\!t\cdot\frac{(-m+k)}{(c_{j}+1+k-1)}\cdot\frac{k+1}{k}\cdot\bar{\alpha}_{j}$
14:   $H^{\prime}_{j}\!\leftarrow\!H^{\prime}_{j}+t$ $\triangleright$ again sign-alternating; compensate
15:   if $|t|<\varepsilon_{\mathrm{rel}}\max(|H^{\prime}_{j}|,\delta_{\mathrm{abs}})$ then break
16:return $B$ and $\{H_{j},H^{\prime}_{j}\}_{j=1}^{d}$ $\triangleright$ used in the backward ratio $H^{\prime}_{j}/H_{j}$
With $H_{j}={}_{2}F_{1}(-m,1;c_{j};\bar{\alpha}_{j})$ and $H^{\prime}_{j}={}_{2}F_{1}(-m+1,2;c_{j}+1;\bar{\alpha}_{j})$, the gradient w.r.t. an $\alpha_{i}$ in bin $S_{j}$ (and zero otherwise) is:
$\frac{\partial B_{\mathrm{H\ddot{o}lder}}}{\partial\alpha_{i}}=-\frac{B_{\mathrm{H\ddot{o}lder}}}{c_{j}}\,\frac{H^{\prime}_{j}}{H_{j}}\cdot\frac{1}{m_{j}},\qquad c_{j}=\frac{q}{b_{j}}+1,$
as derived in §B.4. The same cached $\{H_{j},H^{\prime}_{j}\}$ also feed the zero-aware gap gradients in §B.5 via $\widetilde{{\mathcal{A}}}_{b_{j}}$ (finite sums of $_{2}F_{1}$ with shifted parameters).
Complexity and vectorization.
The forward is $O\!\left(\sum_{j}m\right)=O(dm)$ scalar ops, embarrassingly parallel over bins. The backward reuses the same per-bin computations and adds only $O(dm)$ extra ops for $H^{\prime}_{j}$ and simple scalar multiplications.
Forward objective with zero-aware gap.
[…]
$\mathcal{U}(P)=\underbrace{\prod_{j=1}^{d}\Bigl[{\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)\Bigr]}_{\text{envelope}}^{m_{j}/m}+\ \rho\sum_{j}\underbrace{\frac{m}{2}\,\frac{m_{j}}{m}\,V_{j}^{\omega}\,\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{\alpha}_{j},m\bigr)}_{\text{zero-aware gap}},$
where $V_{j}^{\omega}$ is the within-bin $\omega$-weighted variance ( Proposition 3.8). Both terms reuse the same forward hypergeometric blocks; the second depends only on bin means and the finite differences of the same envelopes.
B.3 Gradient of the envelope for common $\beta$
[…]
${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)\;=\;\frac{1}{q}\;{}_{2}F_{1}\!\Bigl(-m,\,1;\,\tfrac{q}{\beta}+1;\,\bar{\alpha}\Bigr),\qquad q>0,\ \beta>0,\ \bar{\alpha}=\tfrac{1}{m}\sum_{i=1}^{m}\alpha_{i}.$
Since ${\mathcal{H}}_{\beta}$ depends on $\bm{\alpha}$} only through $\bar{\alpha}$, the chain rule gives:
$\frac{\partial{\mathcal{H}}_{\beta}}{\partial\alpha_{i}}\;=\;\frac{\partial{\mathcal{H}}_{\beta}}{\partial\bar{\alpha}}\cdot\frac{\partial\bar{\alpha}}{\partial\alpha_{i}}\;=\;\frac{1}{m}\,\frac{\partial{\mathcal{H}}_{\beta}}{\partial\bar{\alpha}},\qquad i=1,\dots,m.$
Using the standard derivative $\frac{d}{dz}{}_{2}F_{1}(a,b;c;z)=\frac{ab}{c}\;{}_{2}F_{1}(a+1,b+1;c+1;z)$ with $(a,b,c,z)=\bigl(-m,1,\tfrac{q}{\beta}+1,\bar{\alpha}\bigr)$, we obtain:
$\frac{\partial{\mathcal{H}}_{\beta}}{\partial\bar{\alpha}}=\frac{1}{q}\cdot\frac{-m}{\tfrac{q}{\beta}+1}\;{}_{2}F_{1}\!\Bigl(-m+1,\,2;\,\tfrac{q}{\beta}+2;\,\bar{\alpha}\Bigr),$
and hence the per–coordinate gradient:
$\boxed{\;\frac{\partial{\mathcal{H}}_{\beta}(q;\bar{\alpha},m)}{\partial\alpha_{i}}=-\frac{1}{q\bigl(\tfrac{q}{\beta}+1\bigr)}\;{}_{2}F_{1}\!\Bigl(-m+1,\,2;\,\tfrac{q}{\beta}+2;\,\bar{\alpha}\Bigr)=-\frac{\beta}{q(q+\beta)}\;{}_{2}F_{1}\!\Bigl(-m+1,\,2;\,\tfrac{q}{\beta}+2;\,\bar{\alpha}\Bigr),\quad i=1,\dots,m.}$
The gradient is uniform across coordinates because the envelope depends on $\bm{\alpha}$} only via $\bar{\alpha}$. Since m-m is a nonpositive integer, $_{2}F_{1}(-m+1,2;\cdot;\bar{\alpha})$ is a degree-(m1)(m-1) polynomial in $\bar{\alpha}$, enabling stable evaluation via a finite sum or Horner’s rule.
B.4 Gradient of the Hölder envelope for heterogeneous $\beta$
Recall the Hölder envelope (Sec. A.6): with distinct exponents $\{b_{1},\dots,b_{d}\}$, groups $S_{k}\!\vcentcolon=\!\{i:\beta_{i}=b_{k}\}$, sizes $m_{k}\!=\!|S_{k}|$, $m=\sum_{k}m_{k}$, and means $\bar{\alpha}_{k}\!=\!\frac{1}{m_{k}}\sum_{i\in S_{k}}\alpha_{i}$, we defined:
$B_{\mathrm{Holder}}\;=\;\prod_{k=1}^{d}\bigl(B_{k}^{\ast}\bigr)^{m_{k}/m},\qquad B_{k}^{\ast}\;=\;\frac{1}{q}\;{}_{2}F_{1}\!\Bigl(-m,\,1;\,c_{k};\,\bar{\alpha}_{k}\Bigr),\quad c_{k}\vcentcolon=\frac{q}{b_{k}}+1,$
with $q>0$ and $b_{k}>0$. We compute the gradient $\partial B_{\mathrm{Holder}}/\partial\alpha_{i}$ for an index $i\in S_{j}$ (so $\beta_{i}=b_{j}$).
Since $B_{\mathrm{Holder}}$ depends on $\alpha_{i}$ only through $\bar{\alpha}_{j}$,
$\frac{1}{B_{\mathrm{Holder}}}\frac{\partial B_{\mathrm{Holder}}}{\partial\alpha_{i}}\;=\;\frac{m_{j}}{m}\,\frac{1}{B_{j}^{\ast}}\,\frac{\partial B_{j}^{\ast}}{\partial\alpha_{i}},\qquad\frac{\partial B_{j}^{\ast}}{\partial\alpha_{i}}\;=\;\frac{\partial B_{j}^{\ast}}{\partial\bar{\alpha}_{j}}\cdot\frac{\partial\bar{\alpha}_{j}}{\partial\alpha_{i}}\;=\;\frac{1}{m_{j}}\frac{\partial B_{j}^{\ast}}{\partial\bar{\alpha}_{j}}.$
Thus:
$\frac{1}{B_{\mathrm{Holder}}}\frac{\partial B_{\mathrm{Holder}}}{\partial\alpha_{i}}\;=\;\frac{1}{m}\,\frac{1}{B_{j}^{\ast}}\,\frac{\partial B_{j}^{\ast}}{\partial\bar{\alpha}_{j}}.$
Differentiating the hypergeometric.
Using $\frac{d}{dz}{}_{2}F_{1}(a,b;c;z)=\frac{ab}{c}\,{}_{2}F_{1}(a+1,b+1;c+1;z)$ with $(a,b,c,z)=\bigl(-m,1,c_{j},\bar{\alpha}_{j}\bigr)$,
$\frac{\partial B_{j}^{\ast}}{\partial\bar{\alpha}_{j}}\;=\;\frac{1}{q}\cdot\frac{-m}{c_{j}}\;{}_{2}F_{1}\!\Bigl(-m+1,\,2;\,c_{j}+1;\,\bar{\alpha}_{j}\Bigr).$
Combining,
$\frac{1}{B_{\mathrm{Holder}}}\frac{\partial B_{\mathrm{Holder}}}{\partial\alpha_{i}}\;=\;-\frac{1}{q\,c_{j}}\,\frac{{}_{2}F_{1}\!\bigl(-m+1,2;c_{j}+1;\bar{\alpha}_{j}\bigr)}{B_{j}^{\ast}}\;=\;-\frac{1}{c_{j}}\,\frac{{}_{2}F_{1}\!\bigl(-m+1,2;c_{j}+1;\bar{\alpha}_{j}\bigr)}{{}_{2}F_{1}\!\bigl(-m,1;c_{j};\bar{\alpha}_{j}\bigr)},$
since $B_{j}^{\ast}=(1/q)\,{}_{2}F_{1}(-m,1;c_{j};\bar{\alpha}_{j})$. Multiplying by $B_{\mathrm{Holder}}$ yields the per–coordinate gradient (identical for all $i\in S_{j}$, and 0 for $i\notin S_{j}$):
$\boxed{\;\frac{\partial B_{\mathrm{Holder}}}{\partial\alpha_{i}}\;=\;-\frac{B_{\mathrm{Holder}}}{c_{j}}\,\frac{{}_{2}F_{1}\!\bigl(-m+1,2;\,c_{j}+1;\,\bar{\alpha}_{j}\bigr)}{{}_{2}F_{1}\!\bigl(-m,1;\,c_{j};\,\bar{\alpha}_{j}\bigr)},\qquad c_{j}=\frac{q}{b_{j}}+1,\quad i\in S_{j}.\;}$
Equivalent forms.
Let $F_{1}(z)\!\vcentcolon=\!{}_{2}F_{1}(-m,1;c_{j};z)$ and $F_{2}(z)\!\vcentcolon=\!{}_{2}F_{1}(-m+1,2;c_{j}+1;z)$. By the derivative identity, $F_{2}(z)=-\frac{c_{j}}{m}\,\frac{d}{dz}F_{1}(z)$, hence
$\frac{1}{B_{\mathrm{Holder}}}\frac{\partial B_{\mathrm{Holder}}}{\partial\alpha_{i}}\;=\;-\frac{1}{c_{j}}\frac{F_{2}(\bar{\alpha}_{j})}{F_{1}(\bar{\alpha}_{j})}\;=\;\frac{1}{m}\,\frac{d}{dz}\log F_{1}(z)\Big|_{z=\bar{\alpha}_{j}}.$
This gives two numerically equivalent implementations:
(ratio form) $\partial_{\alpha_{i}}\log B_{\mathrm{Holder}}=-\frac{1}{c_{j}}\,\frac{F_{2}(\bar{\alpha}_{j})}{F_{1}(\bar{\alpha}_{j})},$
(log-derivative form) $\partial_{\alpha_{i}}\log B_{\mathrm{Holder}}=\frac{1}{m}\,\frac{d}{dz}\log\!\bigl[{}_{2}F_{1}(-m,1;c_{j};z)\bigr]\Big|_{z=\bar{\alpha}_{j}}.$
Since m-m is a nonpositive integer, both hypergeometric terms truncate:
$_{2}F_{1}(-m,1;c_{j};z)=\sum_{k=0}^{m}\frac{(-m)_{k}(1)_{k}}{(c_{j})_{k}}\frac{z^{k}}{k!}=\sum_{k=0}^{m}(-1)^{k}\binom{m}{k}\frac{z^{k}}{(c_{j})_{k}},$
$_{2}F_{1}(-m+1,2;c_{j}+1;z)=\sum_{k=0}^{m-1}\frac{(-m+1)_{k}(2)_{k}}{(c_{j}+1)_{k}}\frac{z^{k}}{k!}=\sum_{k=0}^{m-1}(-1)^{k}\,\frac{(m-1)!}{(m-1-k)!}\,\frac{k+1}{(c_{j}+1)_{k}}\,z^{k}.$
Thus the ratio in the boxed gradient can be evaluated via stable finite sums (Horner’s rule).
Block structure of the gradient.
For a fixed bin $j$, all coordinates $i\in S_{j}$ share the same partial derivative; for $i\notin S_{j}$ the derivative is zero:
$\frac{\partial B_{\mathrm{Holder}}}{\partial\alpha_{i}}=\begin{cases}-\dfrac{B_{\mathrm{Holder}}}{c_{j}}\,\dfrac{{}_{2}F_{1}(-m+1,2;c_{j}+1;\bar{\alpha}_{j})}{{}_{2}F_{1}(-m,1;c_{j};\bar{\alpha}_{j})},&i\in S_{j},\\[8.0pt] 0,&i\notin S_{j}.\end{cases}$
The log-derivative form is preferred to avoid overflow/underflow when $m$ is large. Note that both $F_{1}$ and $F_{2}$ are nonnegative on $z\in[0,1]$; the gradient is non-positive (increasing any $\alpha_{i}$ weakly decreases the envelope), consistent with the envelope’s monotonicity in $\bar{\alpha}_{j}$. Complexity is O(dm)O(d\,m) per gradient evaluation using the finite sums across bins; computation is easy to parallelize over $j$.
B.5 Gradients of the final objective
For cluster $\ell$ and bin index $j$, let $S_{\ell j}$ be the set of vertices assigned to bin $j$ (with common exponent $b_{j}$), $m_{\ell j}\vcentcolon=|S_{\ell j}|$, $m_{\ell}\vcentcolon=\sum_{j}m_{\ell j}$, and:
$\bar{p}_{\ell j}\;\vcentcolon=\;\frac{1}{m_{\ell j}}\sum_{r\in S_{\ell j}}p_{r\ell},\qquad w_{\ell j}\;\vcentcolon=\;\frac{m_{\ell j}}{m_{\ell}}\quad(\text{Hölder weight}).$
[…]
${\mathcal{H}}_{b_{j}}\bigl(q;\bar{p}_{\ell j},m_{\ell}\bigr)\;=\;\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m_{\ell},1;\tfrac{q}{b_{j}}+1;\bar{p}_{\ell j}\Bigr),$
and the second forward $\beta$–difference and its $\bar{p}$–derivative are ( Proposition 3.8):
${\mathcal{A}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})\;\vcentcolon=\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\,{\mathcal{H}}_{b_{j}}\!\bigl(q+rb_{j};\bar{p}_{\ell j},m_{\ell}\bigr),\qquad\widetilde{{\mathcal{A}}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})\;\vcentcolon=\;\frac{\partial}{\partial\bar{p}_{\ell j}}{\mathcal{A}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell}).$
Zero-aware statistics in a bin.
Fix $i\in S_{\ell j}$ and write
$\omega_{i}\;\vcentcolon=\;\omega(p_{i\ell}),\qquad\Omega_{\ell j}\;\vcentcolon=\sum_{r\in S_{\ell j}}\omega(p_{r\ell}),\qquad S_{2}\;\vcentcolon=\sum_{r\in S_{\ell j}}\omega(p_{r\ell})\,p_{r\ell},$
$\mu\;\vcentcolon=\;\bar{p}^{\omega}_{\ell j}\;=\;S_{2}/\Omega_{\ell j},\qquad V\;\vcentcolon=\;\mathrm{Var}^{\omega}_{\ell j}(p)\;=\;\frac{1}{\Omega_{\ell j}}\sum_{r\in S_{\ell j}}\omega(p_{r\ell})\bigl(p_{r\ell}-\mu\bigr)^{2},$
with the convention $V=0$ if $\Omega_{\ell j}=0$. In the paper we take $\omega(x)=x(1-x)$ so that $\omega^{\prime}_{i}\vcentcolon=\frac{d}{dp}\omega(p_{i\ell})=1-2p_{i\ell}$ (zero-aware and symmetric). The derivatives of the weighted mean and variance are:
$\frac{\partial\mu}{\partial p_{i\ell}}\;=\;\frac{\omega_{i}+\omega^{\prime}_{i}\,(p_{i\ell}-\mu)}{\Omega_{\ell j}},\qquad\frac{\partial V}{\partial p_{i\ell}}\;=\;\frac{1}{\Omega_{\ell j}}\Bigl[\,\omega^{\prime}_{i}\,(p_{i\ell}-\mu)^{2}+2\omega_{i}\,(p_{i\ell}-\mu)\Bigl(1-\frac{\partial\mu}{\partial p_{i\ell}}\Bigr)\,\Bigr]-\frac{V}{\Omega_{\ell j}}\omega^{\prime}_{i}.$ (35)
This is a standard quotient/chain-rule calculation using $\sum_{r\in S_{\ell j}}\omega(p_{r\ell})(p_{r\ell}-\mu)=0$.
Hypergeometric derivatives needed.
Let $q_{r}\vcentcolon=q+rb_{j}$ and $c_{r}\vcentcolon=\tfrac{q_{r}}{b_{j}}+1$. Using $\tfrac{d}{dz}{}_{2}F_{1}(a,b;c;z)=\frac{ab}{c}{}_{2}F_{1}(a+1,b+1;c+1;z)$,
$\frac{\partial}{\partial\bar{p}_{\ell j}}{\mathcal{H}}_{b_{j}}(q_{r};\bar{p}_{\ell j},m_{\ell})$ $=\frac{1}{q_{r}}\cdot\frac{-m_{\ell}}{c_{r}}\;{}_{2}F_{1}\!\Bigl(-m_{\ell}+1,\,2;\,c_{r}+1;\,\bar{p}_{\ell j}\Bigr),$ (36)
$\frac{\partial^{2}}{\partial\bar{p}_{\ell j}^{2}}{\mathcal{H}}_{b_{j}}(q_{r};\bar{p}_{\ell j},m_{\ell})$ $=\frac{1}{q_{r}}\cdot\frac{-m_{\ell}}{c_{r}}\cdot\frac{(-m_{\ell}+1)\cdot 2}{c_{r}+1}\;{}_{2}F_{1}\!\Bigl(-m_{\ell}+2,\,3;\,c_{r}+2;\,\bar{p}_{\ell j}\Bigr).$ (37)
Therefore:
$\widetilde{{\mathcal{A}}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})$ $=\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\,\frac{1}{q_{r}}\cdot\frac{-m_{\ell}}{c_{r}}\;{}_{2}F_{1}\!\Bigl(-m_{\ell}+1,2;c_{r}+1;\bar{p}_{\ell j}\Bigr),$ (38)
$\frac{\partial}{\partial\bar{p}_{\ell j}}\widetilde{{\mathcal{A}}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})$ $=\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\,\frac{1}{q_{r}}\cdot\frac{-m_{\ell}}{c_{r}}\cdot\frac{(-m_{\ell}+1)\cdot 2}{c_{r}+1}\;{}_{2}F_{1}\!\Bigl(-m_{\ell}+2,3;c_{r}+2;\bar{p}_{\ell j}\Bigr).$ (39)
Zero-aware gap term and its gradient.
As stated in the paper, our simple zero-aware upper bound for the AM–GM gap in bin $j$ is:
$\Gamma^{\mathrm{ewa}}_{\ell j}(q)\;\vcentcolon=\;\frac{m_{\ell}}{2}\,w_{\ell j}\;V\;\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{p}_{\ell j},m_{\ell}\bigr),$
so the per–coordinate gradient for $i\in S_{\ell j}$ is:
$\frac{\partial\Gamma^{\mathrm{ewa}}_{\ell j}}{\partial p_{i\ell}}=\frac{m_{\ell}}{2}\,w_{\ell j}\left[\frac{\partial V}{\partial p_{i\ell}}\;\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{p}_{\ell j},m_{\ell}\bigr)\;+\;V\;\frac{\partial\widetilde{{\mathcal{A}}}_{b_{j}}}{\partial\bar{p}_{\ell j}}\cdot\frac{\partial\bar{p}_{\ell j}}{\partial p_{i\ell}}\right],\qquad\frac{\partial\bar{p}_{\ell j}}{\partial p_{i\ell}}=\frac{1}{m_{\ell j}}.$ (40)
Here $\partial V/\partial p_{i\ell}$ and $\partial\mu/\partial p_{i\ell}$ are given by equation 35, while $\widetilde{{\mathcal{A}}}_{b_{j}}$ and its derivative are equation 38–equation 39. For $i\notin S_{\ell j}$, $\partial\Gamma^{\mathrm{ewa}}_{\ell j}/\partial p_{i\ell}=0$.
Envelope term and stick–breaking backward.
The binned Hölder envelope for cluster $\ell$ (Sec. A.6) is:
$B_{\mathrm{Holder},\ell}\;=\;\prod_{k}\Bigl({\mathcal{H}}_{b_{k}}(q;\bar{p}_{\ell k},m_{\ell})\Bigr)^{w_{\ell k}},$
with per–coordinate gradient (for $i\in S_{\ell j}$):
$\frac{\partial B_{\mathrm{Holder},\ell}}{\partial p_{i\ell}}\;=\;-\frac{B_{\mathrm{Holder},\ell}}{c_{j}}\,\frac{{}_{2}F_{1}\!\bigl(-m_{\ell}+1,2;c_{j}+1;\bar{p}_{\ell j}\bigr)}{{}_{2}F_{1}\!\bigl(-m_{\ell},1;c_{j};\bar{p}_{\ell j}\bigr)}\cdot\frac{1}{m_{\ell j}},\qquad c_{j}=\frac{q}{b_{j}}+1,$
obtained by the same log–diff + chain rule used in Sec. B.4. (If the outer objective multiplies the envelope by additional factors—e.g., edge weights $M_{i\ell}(P)$ in the paper—apply product rule and chain through their own Jacobians.)
Putting it together.
Let the final per–cluster contribution be:
$\mathcal{L}_{\ell}(P)\;=\;U_{\ell}(P)\;+\;\rho\sum_{j}\Gamma^{\mathrm{ewa}}_{\ell j}(q),$
where $U_{\ell}$ uses the Hölder envelope (possibly multiplied by problem-specific weights), and $\rho\geq 0$ is the gap regularization. The gradient w.r.t. an entry $p_{i\ell}$ is:
$\frac{\partial\mathcal{L}_{\ell}}{\partial p_{i\ell}}\;=\;\frac{\partial U_{\ell}}{\partial p_{i\ell}}\;+\;\rho\sum_{j:\,i\in S_{\ell j}}\frac{\partial\Gamma^{\mathrm{ewa}}_{\ell j}}{\partial p_{i\ell}},$
with the explicit pieces given in equation 35–equation 40. These feed into the stick–breaking backward pass exactly as in the main text.
If $\Omega_{\ell j}=0$, set $\mu=0$, $V=0$, and $\partial\mu=\partial V=0$; the bin is inactive and contributes no gradient. Because $m_{\ell}$ is a nonpositive integer, all $_{2}F_{1}$ terms truncate to finite polynomials in $\bar{p}_{\ell j}$, enabling stable Horner evaluation for both equation 38 and equation 39. For $\omega(x)=x^{a}$ with $a\in[1,2]$, replace $\omega^{\prime}_{i}$ by $a\,p_{i\ell}^{a-1}$ in equation 35; the rest of the derivation is unchanged.
B.6 A relaxed Hölder envelope via a uniform $c$ and temperature annealing
Setup.
Recall the Hölder envelope (Sec. A.6) for heterogeneous exponents:
$B_{\mathrm{Holder}}\;=\;\prod_{j=1}^{d}\Bigl({\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)\Bigr)^{m_{j}/m},\qquad{\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)=\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;\underbrace{c_{j}}_{=\,q/b_{j}+1};\bar{\alpha}_{j}\Bigr),$
where $b_{j}>0$ is the exponent for bin $j$, $m_{j}=|S_{j}|$, $m=\sum_{j}m_{j}$, and $\bar{\alpha}_{j}=\frac{1}{m_{j}}\sum_{i\in S_{j}}\alpha_{i}$.
Monotonicity in $c$.
For $m\!\in\!\mathbb{N}$ and $z\!\in\![0,1]$, the truncated series:
$_{2}F_{1}(-m,1;c;z)=\sum_{k=0}^{m}\frac{(-m)_{k}(1)_{k}}{(c)_{k}}\frac{z^{k}}{k!}=\sum_{k=0}^{m}(-1)^{k}\binom{m}{k}\frac{z^{k}}{(c)_{k}}$
has nonnegative terms in absolute value and each Pochhammer factor $(c)_{k}=c(c+1)\cdots(c+k-1)$ is strictly increasing in $c$. Hence the whole sum is decreasing in $c$:
$c_{1}\leq c_{2}\ \Longrightarrow\ {}_{2}F_{1}(-m,1;c_{1};z)\ \geq\ {}_{2}F_{1}(-m,1;c_{2};z).$
Within a bin $j$, choose a left–endpoint representative $b_{j}^{\leftarrow}\leq\beta_{i}$ for $i\in S_{j}$ (as in Sec. A.6). Then $c_{j}^{\leftarrow}\!=\!q/b_{j}^{\leftarrow}+1\geq q/\beta_{i}+1$ and, in particular, if $q\!\geq\!b_{j}^{\leftarrow}$ we have $c_{j}^{\leftarrow}\!\geq\!2$. Combining the binwise AM–GM (intra-bin) and Hölder (across bins) steps with the monotonicity equation B.6 yields the relaxed envelope:
${\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)\;=\;\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;c_{j}^{\leftarrow};\bar{\alpha}_{j}\Bigr)\ \leq\ \frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;2;\bar{\alpha}_{j}\Bigr),\quad\text{whenever }c_{j}^{\leftarrow}\geq 2.$
Therefore:
$\boxed{\;B_{\mathrm{Holder}}\ \leq\ \underbrace{\prod_{j=1}^{d}\!\left[\frac{1}{q}\,{}_{2}F_{1}\!\bigl(-m,1;2;\bar{\alpha}_{j}\bigr)\right]^{\!m_{j}/m}}_{\displaystyle\ \vcentcolon=\ B_{\mathrm{relax}(c\!=\!2)}}\qquad(\text{provided }q\geq b_{j}^{\leftarrow}\ \forall j).\;}$
Intuitively, replacing $c_{j}$ by the uniform lower value $2$ (the “largest” case by equation B.6) gives a looser but simpler upper bound. It preserves bin structure through the $\bar{\alpha}_{j}$’s and weights $m_{j}/m$, but removes the explicit $b_{j}$–dependence from the hypergeometric parameter.
Practical simplifications for $c=2$.
Because m-m is a nonpositive integer, $_{2}F_{1}(-m,1;2;z)$ is a degree-$m$ polynomial in $z$ and can be evaluated stably by a finite sum (Horner’s rule):
$_{2}F_{1}(-m,1;2;z)=\sum_{k=0}^{m}(-1)^{k}\,\binom{m}{k}\,\frac{z^{k}}{(2)_{k}}=\sum_{k=0}^{m}(-1)^{k}\,\binom{m}{k}\,\frac{z^{k}}{(k+1)!}.$
Thus:
$B_{\mathrm{relax}(c\!=\!2)}=\prod_{j=1}^{d}\left[\frac{1}{q}\sum_{k=0}^{m}(-1)^{k}\,\binom{m}{k}\,\frac{\bar{\alpha}_{j}^{\,k}}{(k+1)!}\right]^{\!m_{j}/m}.$
This form is handy when one wants to precompute per-bin polynomials in $\bar{\alpha}_{j}$ independent of $b_{j}$.
Temperature–annealed probabilities tighten the zero-aware gap.
Parameterize the assignment probabilities from logits $Z$ at temperature $\tau>0$:
$p_{i\ell}(\tau)=\mathrm{softmax}\!\left(\frac{Z_{i\ell}}{\tau}\right)\quad\text{(multiclass)}\qquad\text{or}\qquad p_{i\ell}(\tau)=\sigma\!\left(\frac{Z_{i\ell}}{\tau}\right)\quad\text{(binary)}.$
As $\tau\downarrow 0$, $p_{i\ell}(\tau)\to\{0,1\}$ elementwise. Our zero-aware gap in bin $j$ uses weights $\omega(x)$ (e.g., $\omega(x)=x(1-x)$ or more generally $\omega(x)=x^{a}$, $a\!\in\![1,2]$), the weighted mean $\mu=\bar{p}_{\ell j}^{\omega}$, and dispersion $V=\mathrm{Var}_{\ell j}^{\omega}(p)$ (Sec. B.5). Two facts hold:
  1. 1.
    Vanishing weights at the extremes. For the choices above, $\omega(0)=\omega(1)=0$ and $0\leq\omega(x)\leq\tfrac{1}{4}$, so for almost-hard assignments $p_{i\ell}(\tau)\in\{0,1\}$ one has $\Omega_{\ell j}(\tau)=\sum_{r\in S_{\ell j}}\omega\bigl(p_{r\ell}(\tau)\bigr)\ \xrightarrow[\tau\downarrow 0]{}\ 0$ and $V(\tau)\xrightarrow[\tau\downarrow 0]{}0$.
  2. 2.
    Zero-aware gap collapses. The (per-bin) gap upper bound
    $\Gamma^{\mathrm{ewa}}_{\ell j}(q)=\frac{m_{\ell}}{2}\,w_{\ell j}\;V\;\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{p}_{\ell j},m_{\ell}\bigr)$
    vanishes as $\tau\downarrow 0$ because $V(\tau)\to 0$ while $\widetilde{{\mathcal{A}}}_{b_{j}}$ stays bounded (finite polynomial in $\bar{p}$). Hence the total objective’s slack from the AM–GM step is driven to zero by temperature annealing.
In contrast, the Hölder envelope terms depend on the bin means $\bar{p}_{\ell j}(\tau)=m_{\ell j}^{-1}\sum_{i\in S_{\ell j}}p_{i\ell}(\tau)$ and thus are insensitive to per-bin dispersion. Annealing shrinks only the gap (and any other dispersion-based penalties), tightening the overall upper bound without altering the envelope’s functional form.
Putting both ideas together.
Combining the relaxed envelope and annealing gives the practical surrogate
$\mathcal{U}_{\mathrm{relax}}(P;\tau)\;=\;\underbrace{\prod_{j=1}^{d}\!\left[\frac{1}{q}\,{}_{2}F_{1}\!\bigl(-m,1;2;\bar{p}_{\ell j}(\tau)\bigr)\right]^{\!w_{\ell j}}}_{\text{uniform-$c$ Hölder envelope}}\;+\;\rho\sum_{j}\underbrace{\frac{m_{\ell}}{2}\,w_{\ell j}\;V_{\ell j}^{\omega}(P(\tau))\;\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{p}_{\ell j}(\tau),m_{\ell}\bigr)}_{\text{zero-aware gap}},$
where $w_{\ell j}=m_{\ell j}/m_{\ell}$ and $\rho\geq 0$. As $\tau\downarrow 0$, $V_{\ell j}^{\omega}(P(\tau))\!\to\!0$ and the gap vanishes, while the envelope is upper-bounded uniformly by the simple $c=2$ hypergeometric polynomial in the bin means.
B.7 Decoupling the qq–dependence
Goal.
For
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\;=\;\int_{0}^{1}\Bigl[\prod_{i=1}^{m}(1-\alpha_{i}+\alpha_{i}t^{\beta_{i}})\Bigr]\,t^{q-1}\,dt,$
we seek bounds of the form
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \underbrace{\Phi(q)}_{\text{$q$ only}}\times\underbrace{\Psi({\bm{\alpha}},{\bm{\beta}},m)}_{\text{independent of $q$}},$
so that tuning/scheduling in $q$ is cleanly separated from model parameters.
Method 1 (variable change): clean decoupling for all $q>0$.
With $t=u^{1/q}$ we have ${\mathcal{I}}(q)=\frac{1}{q}\!\int_{0}^{1}P(u^{1/q})\,du$ and, comparing $u^{\beta_{i}/q}$ to $u^{\beta_{i}}$ on $[0,1]$,
${\mathcal{I}}(q)\ \leq\ \begin{cases}\displaystyle{\mathcal{I}}(1),&q\geq 1,\\[4.0pt] \displaystyle\frac{{\mathcal{I}}(1)}{q},&0<q<1.\end{cases}$
Using the AM–GM envelope at $q=1$, ${\mathcal{I}}(1)\leq H_{1}\vcentcolon={}_{2}F_{1}\!\bigl(-m,1;1/\beta_{\min}+1;\bar{\alpha}\bigr),$ we obtain the decoupled bound
$\boxed{\quad{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \max\!\Bigl(1,\frac{1}{q}\Bigr)\;H_{1}(\bar{\alpha},\beta_{\min},m),\qquad\bar{\alpha}=\tfrac{1}{m}\sum_{i}\alpha_{i}.\quad}$ (40)
Remarks: (i) For $q\geq 1$ a complementary qq–only cap holds, ${\mathcal{I}}(q)\leq\int_{0}^{1}t^{q-1}dt=1/q$, hence ${\mathcal{I}}(q)\leq\min\{H_{1},1/q\}$; equation 40 keeps the qq–factor explicit. (ii) The $\beta_{\min}$ choice is conservative; see the binwise refinement below.
Method 2 (Hölder in $t$): tunable tradeoff with $p>1$.
By Hölder, $\int_{0}^{1}FG\leq\|F\|_{p}\|G\|_{r}$ with $1/p+1/r=1$ and $\|G\|_{r}=\bigl(r(q-1)+1\bigr)^{-1/r}$ for $q>1/p$. Bounding $F(t)=P(t)$ by the qq–free AM–GM envelope at exponent $\beta_{\min}$ yields
${\mathcal{I}}(q)\ \leq\ \underbrace{\Bigl[{}_{2}F_{1}\!\bigl(-mp,1;1/\beta_{\min}+1;\bar{\alpha}\bigr)\Bigr]^{1/p}}_{\Psi({\bm{\alpha}},{\bm{\beta}},m;\,p)}\;\times\;\underbrace{\bigl(r(q-1)+1\bigr)^{-1/r}}_{\Phi(q;\,p)}.$
Example $p=r=2$ (valid for $q>\tfrac{1}{2}$): ${\mathcal{I}}(q)\leq\sqrt{{}_{2}F_{1}(-2m,1;1/\beta_{\min}+1;\bar{\alpha})}\times(2q-1)^{-1/2}$. This family yields potentially tighter decouplings for selected $q$ while keeping $q$ outside the hypergeometric term.
Connection to the Hölder (binning) envelope.
Section A.6 upper-bounds ${\mathcal{I}}(q)$ by a product of per-bin envelopes:
$B_{\mathrm{Holder}}=\prod_{j=1}^{d}\Bigl[{\mathcal{H}}_{b_{j}}\!\bigl(q;\bar{\alpha}_{j},m\bigr)\Bigr]^{m_{j}/m},\qquad{\mathcal{H}}_{b_{j}}(q;\bar{\alpha}_{j},m)=\frac{1}{q}\,{}_{2}F_{1}\!\Bigl(-m,1;\tfrac{q}{b_{j}}+1;\bar{\alpha}_{j}\Bigr).$
Decoupling via the same change of variables at $q=1$ gives a binwise analogue of equation 40:
$\boxed{\quad B_{\mathrm{Holder}}(q)\ \leq\ \max\!\Bigl(1,\frac{1}{q}\Bigr)\;\prod_{j=1}^{d}\Bigl[{}_{2}F_{1}\!\bigl(-m,1;1/b_{j}+1;\bar{\alpha}_{j}\bigr)\Bigr]^{m_{j}/m}.\quad}$ (41)
Thus the qq–dependence is entirely captured by $\max(1,1/q)$; all $\{\alpha_{i},\beta_{i},m\}$ sit inside qq–free per–bin hypergeometric polynomials.
Section B.6 exploited that $_{2}F_{1}(-m,1;c;z)$ is decreasing in $c$ to replace $c_{j}=q/b_{j}+1$ by the uniform $c=2$ (when $q\geq b_{j}^{\leftarrow}$), yielding a simpler polynomial bound. The decoupled versions equation 40–equation 41 play a complementary role: they remove $q$ entirely from the hypergeometric parameter (by evaluating at $q=1$), while §B.6 removes the $b_{j}$–dependence (by fixing $c=2$). Either maneuver produces a fast, qq–factored surrogate.